### v1

# Weighted Aggregation of Conformity Scores for Classification

Abstract
Conformal prediction is a powerful framework for constructing prediction sets with valid coverage guarantees in multi-class classification. However, existing methods often rely on a single score function, which can limit their efficiency and informativeness. We propose a novel approach that combines multiple score functions to improve the performance of conformal predictors by identifying optimal weights that minimize prediction set size. Our theoretical analysis establishes a connection between the weighted score functions and subgraph classes of functions studied in Vapnik-Chervonenkis theory, providing a rigorous mathematical basis for understanding the effectiveness of the proposed method. Experiments demonstrate that our approach consistently outperforms single-score conformal predictors while maintaining valid coverage, offering a principled and data-driven way to enhance the efficiency and practicality of conformal prediction in classification tasks.
1 Introduction
In the era of data-driven decision-making, the ability to quantify uncertainty in machine learning predictions is becoming increasingly important, particularly in high-stakes applications such as medical criminal justice (Berk et al., 2021) and diagnostics (Angelopoulos et al., 2024). While most machine learning methods focus on providing point predictions, supplementing these predictions with valid uncertainty estimates in the form of prediction sets or intervals, is crucial for responsible decision-making.
Conformal prediction (Vovk et al., 2005; Manokhin, 2022) is a powerful framework that provides valid prediction sets with finite-sample coverage guarantees, regardless of the true underlying distribution (Angelopoulos et al., 2023). The key idea is to construct a prediction set for a new test point based on training data, such that the probability of the true label being contained in the prediction set is at least $1-\alpha$. This coverage guarantee holds irrespective of the point prediction algorithm used, making conformal prediction a versatile tool for quantifying uncertainty in machine learning predictions.
Central to conformal prediction is the concept of score functions, which quantify the conformity of a data point to the underlying distribution. In the split conformal prediction framework (Papadopoulos et al., 2002; Lei & Wasserman, 2014; Vovk et al., 2018), the training data is divided into a proper training set and a calibration set. The point prediction algorithm is trained on the proper training set, while the score function is evaluated on the calibration set. The conformal prediction set is then constructed by including all points whose score falls below a certain quantile of the calibration scores, with the quantile determined by the desired coverage level $1-\alpha$.
The choice of score function plays a crucial role in determining the efficiency of the resulting prediction sets, with well-chosen score functions leading to more informative and precise predictions. The flexibility in choosing score functions allows conformal prediction to adapt to the specific characteristics of the data and the underlying distribution. This is an active area of research in conformal prediction, with ongoing efforts to develop score functions that yield the most informative and efficient prediction sets for a given problem setting, including regression (Papadopoulos et al., 2008, 2011; Romano et al., 2019; Kivaranovic et al., 2020; Guan, 2023; Colombo, 2023, 2024) and classification (Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2021; Huang et al., 2024; Luo & Zhou, 2024). Despite the dependence of prediction set efficiency on the choice of score function and the performance of the underlying point prediction algorithm, the coverage guarantee of conformal prediction remains intact, making it a robust tool for providing reliable uncertainty estimates in machine learning predictions.
In this work, we propose a novel approach to optimize a linear combination of conformity scores using efficiency as the objective function. Our goal is to search for optimal weights for such linear combinations, enabling us to fully utilize the advantages of existing score functions. While our approach shares similarities with Izbicki et al. (2019) and Yang & Kuchibhotla (2024), which discuss the construction of the smallest prediction set using kernel density estimators and families of machine learning algorithms, respectively, our method distinguishes itself in three key aspects:
  1. 1.
    First, our approach emphasizes finding an optimal combination of score functions, which is more practical in the current landscape of numerous high-quality and diverse score functions, rather than focusing on training multiple algorithms. By leveraging the strengths of existing score functions, we aim to improve the efficiency of conformal prediction while maintaining its coverage guarantees.
  2. 2.
    Second, we investigate and categorize major data split strategies for determining optimal weights for combining score functions. These strategies include Validity First Conformal Prediction (VFCP) and Efficiency First Conformal Prediction (EFCP) (Yang & Kuchibhotla, 2024), Data Leakage Conformal Prediction (DLCP), and its variant that utilizes all available data (DLCP+). We establish the coverage property for VFCP and quantify the coverage gap for the other three strategies, along with their corresponding efficiency improvements.
  3. 3.
    Third, our method distinguishes itself from selection methods that focus on identifying the single best-performing score function. By aggregating multiple score functions through an optimally chosen weight, our approach can achieve performance levels that surpass those of any individual score function. This ability to harness the collective strengths of multiple score functions sets our method apart and enables it to provide more efficient and informative prediction sets.
Our weighted aggregation approach addresses the challenge of constructing efficient and valid prediction sets in multi-class classification by learning optimal weights for combining different score functions using various data split procedure. This allows for adaptation to the specific characteristics of the dataset and potentially outperforms methods relying on a single score function. We provide a theoretical analysis of our approach, establishing finite-sample guarantees for coverage and width and deriving oracle inequalities to quantify its performance relative to an “optimal” aggregation. Through extensive experiments, we demonstrate the superior performance of our weighted aggregation approach compared to single-score methods, highlighting its potential for improving the efficiency of conformal prediction in classification tasks.
The rest of the paper is organized as follows. In Section 2, we present our weighted aggregation approach and the various data split strategies. Section 3 contains our theoretical analysis, establishing coverage guarantees and deriving the expected prediction set sizes. In Section 4, we discuss our experimental results, demonstrating the effectiveness of our approach. Related works are presented in Section 5. Finally, we conclude the paper in Section 6 and discuss future research directions.
2 Weighted Aggregation of Score Functions
2.1 Basic Conformal Prediction for Classification
We start by assuming that a classification algorithm provides $\widehat{p}_{y}(x)$, which approximates $P(Y=y|X=x)$. While our method and theoretical analysis do not depend on the accuracy of this approximation, it is beneficial to assume that higher values of $\widehat{p}_{y}(x)$ indicate a greater likelihood of sample $x$ having label $y$. We consider this training procedure to be performed on a separate dataset, ensuring that $\widehat{p}(y|x)$ is independent of the dataset used in this paper.
Let us first consider a single non-conformity score function $s(x,y)$. This function is defined such that higher values of $s(x,y)$ indicate a higher priority of believing $x$ has label $y$. A common choice for the non-conformity score is $s(x,y)=\widehat{p}_{y}(x)$ Sadinle et al. (2019). In this context, conformal prediction for classification problems can be described by the following algorithm. To summarize, we first find a threshold in a set of labeled data, so that $s(x_{i},y_{i})\geq Q_{1-\alpha}$ holds for at least $1-\alpha$ proportion in the set $\mathcal{I}_{\textrm{train}}$. Then we use this threshold to define the prediction set for $x_{i}$ in the test set, which is the upper level set of the function $y\mapsto s(x_{i},y)$.
Algorithm 1 Split Conformal Prediction
1:Data $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$, $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$, pre-determined coverage probability $1-\alpha$
2:A prediction set $\widehat{C}_{\alpha}(x_{i}),i\in\mathcal{I}_{\textrm{test}}$
3:Randomly split $\mathcal{I}_{\textrm{train}}$ into $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
4:Train a model $\widehat{p}(x)$ on $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$.
5:$Q_{1-\alpha}\leftarrow\lceil(1+|\mathcal{I}_{2}|)(1-\alpha)\rceil$-th smallest non-conformity score $s(x_{i},y_{i})$ for $i\in\mathcal{I}_{2}$.
6:$\widehat{C}_{1-\alpha}(x_{i},Q_{1-\alpha})\leftarrow\{y\in[K]:s(x_{i},y)\geq Q_{1-\alpha}\}$ for $i\in\mathcal{I}_{\textrm{test}}$.
This algorithm constructs prediction sets $\widehat{C}_{1-\alpha}(x_{i})$ for each $i\in\mathcal{I}_{\textrm{test}}$, based on the non-conformity scores and a threshold determined by the desired coverage probability $1-\alpha$. The algorithm splits the training data into two subsets: $\mathcal{I}_{1}$ for calculating the non-conformity scores, and $\mathcal{I}_{2}$ for calibration. The threshold $Q_{1-\alpha}$ is chosen to ensure the desired coverage probability.
A key property of this method is that, under the assumption of exchangeability of the data points (a weaker assumption than i.i.d.), it guarantees that the resulting prediction sets will contain the true label with probability at least $1-\alpha$. This property holds regardless of the accuracy of the underlying classification algorithm, making conformal prediction a powerful tool for uncertainty quantification.
2.2 Example of Score Functions for Classifications
In the context of conformal prediction, the choice of score function $s(x,y)$ is crucial as it directly impacts the resulting prediction sets. These score functions serve as non-conformity measures, quantifying how well a label fits a given input. Let $K$ denote the number of classes in our classification problem.
Various score functions have been proposed and their effectiveness demonstrated through extensive experimentation on diverse datasets. Three typical choices of score functions are:
  1. 1.
    Threshold (THR) (Sadinle et al., 2019): $s_{1}(x,y)=\widehat{p}_{y}(x)$.
  2. 2.
    Adaptive Prediction Sets (APS) (Romano et al., 2020): $s_{2}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\}$.
  3. 3.
    Rank-based (RANK) (Luo & Zhou, 2024): $s_{3}(x,y)=|\{y^{\prime}\in[K]:\widehat{p}_{y^{\prime}}(x)<\widehat{p}_{y}(x)\}|$.
We can easily plug the APS and RANK score functions into Algorithm 1 to create prediction sets. While the original publications introducing these score functions presented additional complexities to guarantee precise $1-\alpha$ expected coverage, these adjustments do not significantly impact the presentation of our proposed approach and are therefore omitted for clarity. The reasoning behind these scores will not affect how we combine them or our theoretical analysis. In the following sections, we will demonstrate how these score functions can be effectively utilized in our proposed approach without the need for additional complexity to ensure exact coverage. A detailed introduction and discussion of these score functions, along with their adjustments, will be provided in Section 4.
2.3 Aggregation of Score Functions
In Yang & Kuchibhotla (2024), the authors propose aggregating score functions for conformal prediction. Their approach selects the score function that yields the smallest average prediction set size among the available score functions. They also introduce two data splitting methods: Efficiency First Conformal Prediction (EFCP) and Validity First Conformal Prediction (VFCP).
Given a vector of score functions $s(x,y)=(s_{1}(x,y),\dots,s_{d}(x,y))^{\top}$, instead of simply choosing the score that provides the smallest prediction set on a calibration set, we propose assigning weights to the scores and defining a new weighted score function:
$\langle\bm{w},s(x,y)\rangle=\sum_{j=1}^{d}\bm{w}_{j}s_{j}(x,y).$ (1)
By leveraging different scores to create a completely new score function, we expect the weighted score to outperform any individual score. This approach allows for more flexibility in combining the strengths of various score functions, potentially leading to improved prediction set efficiency.
2.4 Algorithm for Determining the Weight and the Threshold
In this section, we present a detailed procedure for determining the optimal weight vector $\bm{w}$ in the context of the weighted score approach introduced in Eq. (1). This procedure extends Algorithm 1 by incorporating an additional step to minimize the expected size of the prediction set. The following steps outline the process:
  1. 1.
    Extract $\mathcal{I}_{1},\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ from the dataset. In the meta version of the algorithm, these sets can be arbitrary, provided they are sufficiently large. However, for clarity and ease of understanding for first-time readers, it is helpful to assume that $\mathcal{I}_{1}=\mathcal{I}_{2}$ and that $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are partitions of $\mathcal{I}_{\textrm{train}}$.
  2. 2.
    Define a finite subset $\mathcal{W}\subseteq\mathbb{R}^{d}$ for candidates of $\bm{w}$. Theoretically, $\mathcal{W}$ can be an arbitrary subset of $\mathbb{R}^{d}$. Given that the scores are all non-conformity scores, it is sometimes desirable for the weights to be non-negative. In this case, letting $\mathcal{W}$ be the grid of the simplex $\Delta(d)\subseteq\mathbb{R}^{d}$ is a typical choice. The compactness of $\Delta(d)$ can make the algorithm computationally feasible.
  3. 3.
    Estimate the threshold for every $\bm{w}\in\mathcal{W}$ on $\mathcal{I}_{1}$. In the set $\mathcal{I}_{1}$, calculate the threshold for every $\bm{w}\in\mathcal{W}$:
    $Q_{1-\alpha}^{(1)}(\bm{w})=\lceil(1+|\mathcal{I}_{1}|)(1-\alpha)\rceil\text{-th smallest value of }\{\langle\bm{w},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{1}\}.$ (2)
    This corresponds to Step 5 in Algorithm 1.
  4. 4.
    Determine the prediction set for data in $\mathcal{I}_{2}$. For a given $\bm{w}$ and threshold $Q_{1-\alpha}^{(1)}(\bm{w})$, define the prediction set for each sample $x_{i},i\in\mathcal{I}_{2}$ as:
    $\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))=\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(1)}(\bm{w})\}\quad\text{for}\quad i\in\mathcal{I}_{2}.$ (3)
    This corresponds to Step 6 in Algorithm 1. However, we are not done yet, since these are the prediction sets for all $\bm{w}\in\mathcal{W}$, and our goal is to minimize the prediction set size. This intuitively leads to the next step.
  5. 5.
    Find the optimal weight vector $\widehat{\bm{w}}$ that minimizes the average prediction set size:
    $\widehat{\bm{w}}\in\arg\min_{\bm{w}\in\mathcal{W}}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))|.$ (4)
  6. 6.
    If $\mathcal{I}_{3}\neq\mathcal{I}_{1}$, then $\mathcal{I}_{3}$ is treated as the calibration set in Step 5 in Algorithm 1, and we obtain:
    $Q_{1-\alpha}^{(2)}=\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil\text{-th smallest value of }\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{3}\}.$ (5)
    If $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are disjoint, $\widehat{\bm{w}}$ and $Q_{1-\alpha}^{(2)}$ are independent. If $\mathcal{I}_{1}=\mathcal{I}_{3}$, then this quantile has been computed in Eq. (2), i.e., $Q_{1-\alpha}^{(2)}=Q_{1-\alpha}^{(1)}(\widehat{\bm{w}})$.
  7. 7.
    The final output is the confidence set:
    $\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})=\{y\in[K]:\langle\widehat{\bm{w}},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(2)}\}\quad\text{for}\quad i\in\mathcal{I}_{\textrm{test}}.$ (6)
The following two key observations of this procedure:
  • (a)
    Once the optimal weight vector $\widehat{\bm{w}}$ has been determined, the subsequent steps of the calibration procedure, as described in Eq. (5), and the construction of the prediction set, as outlined in Eq. (6), are identical to Step 5 and Step 6 of Algorithm 1. The additional steps preceding these are designed to identify an appropriate weight vector $\bm{w}$ that optimizes the performance of the algorithm.
  • (b)
    The operation in Eq. (2) and Eq. (5) are the steps for finding quantiles, and Eq. (3) and Eq. (6) are the steps for defining prediction sets. For finding the quantiles, we need the labels of the samples, while for defining the prediction set, only the feature $x$ of the sample is required. In other words, it is possible to include the test set in Eq. (3) and Eq. (6).
In the following algorithm, we present a formalized version of the procedure outlined above.
Algorithm 2 Optimal Weight Selection and Prediction Set Construction
1:$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$: Labeled training data; $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$: Unlabeled test data; $\alpha\in(0,1)$: Pre-defined significance level; $\mathcal{W}\subseteq\mathbb{R}^{d}$: Set of possible weight vectors, where $d$ is the number of score functions.
2:$\big{(}C_{1-\alpha}(X_{i})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$: Sequence of prediction sets for test data with anticipated $1-\alpha$ coverage
3:Decide $\mathcal{I}_{1}\subseteq\mathcal{I}_{\textrm{train}}$, $\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$, and $\mathcal{I}_{3}\subseteq\mathcal{I}_{\textrm{train}}$ for Step 1, 2, and 3, respectively
4:Step 1: Grid Search over Weights
5:for $\bm{w}\in\mathcal{W}$ do
6:     $Q_{1-\alpha}^{(1)}(\bm{w})\leftarrow$ Calibration($\bm{w},\mathcal{I}_{1},\alpha$).
7:end for
8:Step 2: Optimal Weight Selection
9:for $\bm{w}\in\mathcal{W}$ do
10:     $\big{(}\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}\leftarrow$ Evaluation($\bm{w},\mathcal{I}_{2},Q_{1-\alpha}^{(1)}(\bm{w})$).
11:     Compute average prediction set size $S{(\bm{w})}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w})\big{)}|$.
12:end for
13:$\widehat{\bm{w}}\leftarrow\arg\min_{\bm{w}\in\mathcal{W}}S{(\bm{w})}$.
14:Step 3: Test-time Prediction Set Construction
15:$Q_{1-\alpha}^{(2)}\leftarrow$ Calibration($\widehat{\bm{w}},\mathcal{I}_{3},\alpha$).
16:$\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}\leftarrow$ Evaluation($\widehat{\bm{w}},\mathcal{I}_{\textrm{test}},\mathcal{I}_{3}$).
17:return $\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$
Function 1 Calibration
1:function Calibration($\bm{w},\mathcal{I}\subseteq\mathcal{I}_{\textrm{train}},\alpha$)
2:     for $i\in\mathcal{I}_{\textrm{calib}}$ do
3:         Compute score $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
4:     end for
5:     $Q_{1-\alpha}\leftarrow$ $\lceil(1+|\mathcal{I}|)(1-\alpha)\rceil$-th smallest score $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
6:     return $Q_{1-\alpha}$.
7:end function
Function 2 Evaluation
1:function Evaluation($\bm{w},\mathcal{I}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}},Q_{1-\alpha}$)
2:     for $i\in\mathcal{I}$ do
3:         for $y\in[K]$ do
4:              Compute non-conformity score $\langle\bm{w},s(x_{i},y)\rangle$ using $\bm{w}$ and $(x_{i},y)$
5:         end for
6:         $\widehat{C}_{1-\alpha}(x_{i})\leftarrow\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}\}$
7:     end for
8:     return $\big{(}\widehat{C}_{1-\alpha}(x_{i})\big{)}_{i\in\mathcal{I}}$
9:end function
For $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$, we present an overview of the data split framework in Figure 1. For what follows, we introduce the following four choices.
  • (a)
    For Validity First Conformal Prediction (VFCP) (Yang & Kuchibhotla, 2024), $\mathcal{I}_{1}=\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}\setminus\mathcal{I}_{1}$. This method divides $\mathcal{I}_{\textrm{train}}$ into two partitions, using less data to determine $\widehat{\bm{w}}$. However, the marginal coverage probability is guaranteed to be at least $1-\alpha$ under the assumption of exchangeability of samples in $\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$. The name “validity first” comes from this fact.
  • (b)
    For Efficiency First Conformal Prediction (EFCP) (Yang & Kuchibhotla, 2024), $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$. This method uses all training data to determine $\widehat{\bm{w}}$, resulting in a more accurate estimation of the optimal $\bm{w}$. However, stronger assumptions are required for the coverage guarantee. The name “efficiency first” comes from this fact.
  • (c)
    For Data Leakage Conformal Prediction (DLCP), $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. Since $\mathcal{I}_{2}$ can be unlabeled data, we can minimize the prediction set on $\mathcal{I}_{\textrm{test}}$ in the step of finding $\widehat{\bm{w}}$ in Eq. (4) by letting $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. We call this a “data leakage” method because we use the test data in the training procedure of $\widehat{\bm{w}}$.
  • (d)
    For another version of Data Leakage Conformal Prediction (DLCP+), $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. If data leakage is allowed, then using the same idea as EFCP, we can use all available data in each step. Therefore, we can include $\mathcal{I}_{\textrm{test}}$ in $\mathcal{I}_{2}$ to maximize the size of samples in the step of finding $\widehat{\bm{w}}$ in Eq. (4).
$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{3}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{2}}$ $Q_{1-\alpha}^{(1)}(\bm{w})$ by Eq. (2) $\big{(}\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$ by Eq. (3) $\widehat{\bm{w}}$ by Eq. (4) $Q_{1-\alpha}^{(2)}$ by Eq. (5) $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$ $\left(\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\right)_{i\in\mathcal{I}_{\textrm{test}}}$ by Eq. (6)
Figure 1: This example shows an overall framework of data splitting for $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$. The complete procedure is presented in Algorithm 2. Shortly speaking, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are used in Step 1 and 2 to select optimal weight $\widehat{\bm{w}}$, $\mathcal{I}_{3}$ is used in Step 3 as calibration set to construct prediction sets for $\mathcal{I}_{\textrm{test}}$. Under this framework, we illustrate four options, VFCP, EFCP, DLCP, and DLCP+, and discuss their coverage and size property in Section 3 as well as empirical results in Section 4.
3 Theoretical Analysis
The theoretical analysis of our method are based on the subgraph classes in Vapnik–Chervonenkis (VC) theory. This theoretical details and the proofs of lemma and theorems will be left in Section A and B in the Appendix. In this section, we will present the key result for our proposed algorithm. The output of the algorithm analyzed in two aspect:
  1. 1.
    Can the prediction set achieve $1-\alpha$ coverage?
  2. 2.
    Can the prediction set achieve the minimal size? The minimal size will be defined as below.
$\bm{w}^{*}$ is defined as the population level of $\widehat{\bm{w}}.$
$\bm{w}^{*}=\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}[|\widehat{C}_{1-\alpha}(X,\bm{w},q)|]\quad{s.t.}\quad\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha.$
Equivalently, we can define
$Q_{1-\alpha}(\bm{w})=\sup\{q\in\mathbb{R}:P(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha\},$
and using this definition, we define
$\bm{w}^{*}=\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}[|\widehat{C}_{1-\alpha}(X,\bm{w},Q_{1-\alpha}(\bm{w})|]$
Let us define $q^{*}=Q_{1-\alpha}(\bm{w}^{*})$. Since $\bm{w}^{*}$ and $q^{*}$ are deterministic once $\alpha$ and the true distribution are given, We can briefly denote the optimal prediction set for $x$ by
$\widehat{C}^{*}_{1-\alpha}(x):=\widehat{C}_{1-\alpha}(x,\bm{w}^{*},q^{*}).$
From a theoretical perspective, the expected size of the prediction set (at the population level) obtained using $\widehat{\bm{w}}$ cannot be smaller than the one provided by $\bm{w}^{*}$. In the subsequent analysis, we will investigate the difference between the prediction set sizes yielded by $\widehat{\bm{w}}$ and $\bm{w}^{*}$. To facilitate this analysis, we will introduce and rely on the following events. We define
$\Omega(\mathcal{I},\eta)=\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq q\}-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq q\}]\right|\leq\eta$ (7)
and
$\Gamma(\mathcal{I},\xi)=\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(X_{i},y)\rangle\geq q\}-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,y)\rangle\geq q\}]\right|\leq\xi.$ (8)
$\Omega(\mathcal{I},\eta)$ is about the uniform concentration of the coverage, and $\Gamma(\mathcal{I},\xi)$ is about the uniform concentration of the prediction set size. This is because by the definition of $\widehat{C}_{1-\alpha}$ in Eq. (2) or Eq. (5),
$\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\mathbf{1}\big{\{}y\in\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{\}}\quad{\text{and}}\quad\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\big{|}\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{|}.$
The following lemma shows that these two events hold with high probability.
Lemma 1.
Suppose the samples in $\mathcal{I}$ are i.i.d., then
$\mathbb{P}\left(\Omega\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (9)
and
$\mathbb{P}\left(\Gamma\left(\mathcal{I},8K\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+K\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (10)
This lemma says that $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with high probability if $\eta,\xi\gtrsim\sqrt{\frac{d\log|\mathcal{I}|}{|\mathcal{I}|}}$. Now we are ready to introduce the statistical guarantees of the methods.
3.1 Consistency of VFCP
The statistical guarantees of VFCP requires a few additional assumptions, e.g., the continuity between the quantile and the prediction set. To get rid off such assumptions and present the result in a neater way, we will only prove a modified version of VFCP. We will change $\alpha$ in Eq. (2) to a slightly smaller $\alpha^{\prime}$. The condition for $\alpha^{\prime}$ will be specified in the theorem.
Theorem 1.
Let $\widehat{C}_{1-\alpha}^{\text{VFCP}}(x)$ be the output of Algorithm 2 with the setting of VFCP, and modified as mention above. Suppose the samples in $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable, then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\mid\widehat{\bm{w}}\right)\geq 1-\alpha.$
Moreover, if $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, and $\alpha^{\prime}+\eta_{1}+\eta_{3}\leq\alpha$, then for any $X$ in the test set,
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\big{|}\mid\widehat{\bm{w}},\widehat{Q}_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\ \right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{2},$
where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha^{\prime})\rceil-\eta_{3}.$
In the setting of EFCP, $\mathcal{I}_{1}=\mathcal{I}_{2}$. To emphasize their roles in the algorithm, we use different notation in the theorem. $\widehat{\bm{w}}$ and $\widehat{Q}_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ depend on the dataset $\mathcal{I}_{\textrm{train}}.$ As long as $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, then the coverage and the expected prediction set size satisfies the inequalities in the theorem.
3.2 Consistency of EFCP
In the setting of EFCP, $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$. We denote this set by $\mathcal{I}_{\textrm{train}}$.
Theorem 2.
Let $\widehat{C}_{1-\alpha}^{\text{EFCP}}(x,\widehat{\bm{w}})$ be the output of Algorithm 2 with the setting of EFCP. Suppose the data in $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ are independent and $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ is satisfied. Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{EFCP}}(X;\widehat{\bm{w}})\mid\widehat{\bm{w}}\ \right)\geq 1-\alpha_{1}\quad\text{where}\quad 1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
Moreover, if $\Gamma(\mathcal{I}_{\textrm{train}},\xi_{\text{train}})$ is satisfied, then for any $X$ in the test set,
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{EFCP}}(X,\widehat{\bm{w}})\big{|}\mid\widehat{\bm{w}}\ \right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{\text{train}}.$
3.3 Consistency of DLCP
In the setting of DLCP, $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. We will denote this by $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ respectively. In this case, the test set and $\widehat{\bm{w}}$ become dependent. We will present the result under a high probability event about the prediction set size.
Theorem 3.
Let $\widehat{C}_{1-\alpha}^{\text{DLCP}}(x))$ be the output of Algorithm 2 with the setting of DLCP. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})\}\geq 1-\alpha_{1}-\eta_{\text{test}}\quad\text{where}\quad 1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
In addition, if $\Gamma(\mathcal{I}_{\textrm{test}},\xi_{\text{test}})$ holds, then
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}|\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})|\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+\xi_{\text{test}}.$
3.4 Consistency of DLCP+
Similar as the DLCP setting, $\widehat{\bm{w}}$ depends on the test set, so the result of prediction set size is presented in a similar way.
Theorem 4.
Let $\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x,\widehat{\bm{w}}))$ be the output of Algorithm 2 with the setting of DLCP+. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x_{i},\widehat{\bm{w}})\}\geq 1-\alpha^{\prime}-\eta_{\text{test}}\quad\text{where}\quad 1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
In addition, if $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for $X$ in the test set,
$\mathbb{E}[|\widehat{C}_{1-\alpha}^{\text{DLCP+}}(X)|]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha^{\prime}}(X)\big{|}\right]+2\xi_{2}.$
We note that in the setting of DLCP+, $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. We keep use $\mathcal{I}_{2}$ in the theorem to avoid long notation.
4 Experiments
We conducted an experiment to compare the performance of various score functions and their weighted combinations for prediction set evaluation. We focused on three powerful score functions: THR (Sadinle et al., 2019), APS (Romano et al., 2020), and RANK (Luo & Zhou, 2024). Throughout the experiment, we assumed that a pretrained classifier $\widehat{p}$ was available, and we focused on the split of the dataset unseen during the training of the pretrained classifier [^1] The unseen dataset was further divided into $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$. The naming for $\mathcal{I}_{\textrm{train}}$ is not to be confused with the dataset for training. .
The three basic score functions used in our experiment were:
Least Ambiguous Set Values Classifier (THR) (Sadinle et al., 2019).
The score function of THR is defined as:
$s_{\text{THR}}(x,y)=\widehat{p}_{y}(x).$
This score function is straightforward: it assigns a higher score to labels with a higher estimated probability. The prediction set includes labels with the highest estimated probabilities. This is the score function we have to include because if $\widehat{p}$ is the true posterior probability, then the score function itself can achieve the smallest expected prediction set size.
Adaptive Prediction Set (APS) (Romano et al., 2020).
The score function of APS is defined as:
$s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\},$
where $\mathbf{1}\{\cdot\}$ is the indicator function. This score function can be interpreted as the complement of the $p$-value for label $k$. It measures the sum of the estimated probabilities of all labels that have the same or a smaller estimated probability than label $k$.
Rank-based Score Function (RANK) (Luo & Zhou, 2024).
The score function of RANK is defined as:
$s_{\text{RANK}}(x,y)=\frac{|\{k^{\prime}\in[K]:\widehat{p}_{k^{\prime}}(x)<\widehat{p}_{k}(x)\}|}{K-1},$
This score function assigns a score based on the rank of the estimated probability $\widehat{p}_{y}(x)$ among all the estimated probabilities for input $x$. The rank is divided by $K-1$ so that the range of the score is from 0 to 1. The prediction set gives higher priority to labels with larger ranks.
The three score functions, THR, APS, and RANK, employ different strategies for assigning scores based on the estimated probabilities. THR directly utilizes the estimated probabilities as scores. APS, on the other hand, considers the cumulative probability of labels that have the same or lower estimated probabilities compared to the label of interest. RANK, in contrast, focuses on the relative ranking of the estimated probabilities among all possible labels. It is important to note that all these score functions preserve the order of the labels, which means that, for fixed $x$, the order of labels based on their estimated probabilities, $\widehat{p}_{y}(x)$, remains the same when ranked according to their scores, $s(x,y)$.
Figure 2: Boxplot comparison of different score functions at a significance level of $\alpha=0.01$. Our weighted combination method achieves the guaranteed coverage of 99% while maintaining the smallest prediction set size among all the compared score functions.
Experiments were conducted on the CIFAR-100 dataset (Alex, 2009) using a pretrained RepVGG model (Ding et al., 2021). Our experiments focus on this dataset because of the following facts we found in our experiments.
  1. 1.
    This dataset has a well pretrained model, RepVGG (Ding et al., 2021) as our benchmark model.
  2. 2.
    This is one of the most challenging classification tasks in computer vision because of its large number of classes.
  3. 3.
    There is no existing score functions proposed in earlier work that obviously outperforms all others.
Given these three aspects, it will be easier for us to present the effectiveness of the weighted score.
The 10000 testing images, which were not used during the pretraining of the model, were used as the $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ sets. The experiments were performed for different significance levels $\alpha$ ranging from 0.01 to 0.05 and two different ratios of $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$, namely 80:20 and 99:1. For each combination of $\alpha$ and data split ratio, 100 runs with different index splits were conducted to ensure robustness.
The main focus of the experiments was to compare the performance of the proposed weighted score function with three constituent base score functions: APS, THR, and RANK. Additionally, the proposed method was compared with two competitive baseline score functions, RAPS (Angelopoulos et al., 2021) and SAPS (Huang et al., 2024).
Figure 2 displays the boxplot of the coverage and size of our proposed method alongside five other score functions at a significance level of $\alpha=0.01$. To further analyze the performance across different significance levels, Figure 3 presents a plot of the size against coverage for the various score functions, with $\alpha$ values ranging from 0.01 to 0.05. To ensure a fair comparison and guarantee the coverage property, the first data split method, VFCP, was employed for all methods. Our method achieves the guaranteed coverage with the smallest prediction set size among all compared score functions, demonstrating its ability to optimize efficiency while maintaining the desired coverage level.
Figure 3: Comparison of size vs. coverage for different score functions as well as our proposed method across $\alpha$ values ranging from 0.01 to 0.05. The results indicate that our weighted combination method (shown in red) consistently outperforms the three constituent basic score functions (APS, THR, RANK) as well as other competitive baseline score functions (RAPS, SAPS) in terms of achieving coverage with smaller prediction set sizes.
To investigate the impact of various data split methods (Section 2.4), we further compare the performance of our weighted score function under different split approaches. Additionally, we include the results of the five individual score functions using the VFCP split method. Table 1 offers a comprehensive overview of the coverage and size results at significance levels of $\alpha=0.01$ and 0.05. These findings demonstrate the effectiveness of the different split strategies.
$\alpha=0.01$ $\alpha=0.05$
Method Coverage Size Coverage Size
Weighted (VFCP) 0.990 (0.003) 13.782 (1.114) 0.950 (0.005) 3.890 (0.266)
Weighted (EFCP) 0.989 (0.003) 13.306 (0.464) 0.949 (0.005) 3.754 (0.096)
Weighted (DLCP) 0.989 (0.003) 13.298 (0.461) 0.949 (0.005) 3.752 (0.097)
Weighted (DLCP+) 0.989 (0.003) 13.299 (0.459) 0.949 (0.005) 3.753 (0.097)
APS 0.990 (0.003) 40.217 (1.786) 0.949 (0.006) 19.545 (1.022)
THR 0.989 (0.003) 26.949 (2.198) 0.949 (0.006) 4.519 (0.381)
RANK 0.990 (0.003) 27.161 (2.421) 0.950 (0.006) 6.298 (0.395)
RAPS 0.990 (0.003) 26.403 (2.505) 0.950 (0.005) 5.581 (0.263)
SAPS 0.990 (0.003) 25.096 (2.298) 0.950 (0.006) 5.036 (0.302)
Table 1: Coverage and size of different methods for $\alpha=0.01$ and $\alpha=0.05$ on CIFAR-100 dataset. Results are shown as mean (standard deviation).
(a) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 80:20. EFCP, DLCP, and DLCP+ exhibit similar size results, but DLCP has the smallest coverage and the largest gap from the desired coverage level of $0.951-\alpha=$ 0.95. VFCP attains the desired coverage at the cost of having the largest prediction set size.
(b) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 99:1. DLCP achieves the smallest size among the compared methods. VFCP attains the desired coverage at the cost of having the largest prediction set size.
Figure 4: Comparison of coverage and size for different data split methods at $\alpha$=0.05.
5 Related Work
We would like to underscore the connections between our work and existing studies on conformal prediction, particularly those that delve into model aggregation and calibration. Regarding model aggregation, Yang & Kuchibhotla (2024) introduce two selection algorithms aimed at minimizing the width of the prediction interval by aggregating and selecting from regression estimators. Stutz et al. (2023) further extends conformal prediction to scenarios where the ground truth label is ambiguous, employing predictions from multiple labels to construct valid prediction sets.
In the context of model calibration, temperature scaling (Guo et al., 2017) is utilized as a post hoc method to adjust classifier confidence levels. Additionally, Rahimi et al. (2020) endeavors to calibrate predicted probabilities from deep learning models while maintaining the rank order of these probabilities. Dabah & Tirer (2024); Xi et al. (2024) empirically demonstrate that applying post hoc calibration prior to conformal prediction can reduce model overconfidence but tends to increase the size of the prediction interval. Lin et al. (2021) addresses calibration in regression settings, leveraging kernel regression and localized conformal prediction based on a pretrained neural network mean estimator to create valid and discriminative prediction intervals.
However, the original split conformal method (Papadopoulos et al., 2002) only utilizes a portion of the data for prediction, leading to reduced data efficiency. To address this limitation, several approaches have been proposed to enhance data utilization while maintaining computational efficiency.
Aggregation methods have emerged as a popular approach to improve data utilization as well as prediction efficiency in conformal prediction. Carlsson et al. (2014) extended inductive conformal prediction (ICP) by aggregating predictions from multiple ICPs, reducing variance and improving efficiency, albeit more conservatively than ICP. The cross-conformal prediction Vovk (2015); Vovk et al. (2018) presents a hybrid of inductive conformal prediction and cross-validation, offers computational efficiency comparable to ICP while producing valid predictions empirically. The validity of aggregated conformal predictors is not fully understood and requires additional assumptions; a revised definition is proposed Linusson et al. (2017) to achieve approximate validity conditional on the underlying learning algorithm’s properties. Johansson et al. (2014) demonstrated that random forests as the underlying model for conformal regression outperform others in terms of efficiency, using out-of-bag estimates instead of a separate calibration set. BARBER et al. (2021) introduced the jackknife+ method, which employs leave-one-out cross-validation to compute nonconformity scores for each example in the calibration set, enabling the use of all available data for both training and calibration. Kim et al. (2020) proposed a similar approach using K-fold cross-validation, where the model is trained on multiple folds of the data, and the predictions from each fold are combined to create a valid prediction set. These methods strike a balance between data efficiency and computational complexity by reusing data for both training and calibration. Recently, conformal prediction has been extended to handle distribution shifts Tibshirani et al. (2019); Gibbs et al. (2023); Prinster et al. (2024), relaxing the assumption of exchangeability. This advancement allows for the application of conformal prediction to a wider range of real-world scenarios, such as data with dependencies Chernozhukov et al. (2021); Luo et al. (2023); Nettasinghe et al. (2023); Luo & Colombo (2024).
It is worth noting that there is an inherent connection between model aggregation and our approach of weighting non-conformity scores, even though our method does not directly address the scenario of combining multiple models to find an optimal aggregation.
To illustrate this connection, consider a scenario with three classifiers, including an oracle classifier making perfect predictions. A model aggregation method such as boosting (Mukherjee & Schapire, 2010) would assign a weight of 1 to the oracle classifier, disregarding the others. Similarly, when weighting non-conformity scores, even without a perfect classifier, the resulting conformal prediction sets will be equivalent to applying the THR score to the oracle classifier. The optimally chosen weighted non-conformity score can identify and prioritize the most informative and reliable scores, mimicking an oracle classifier. This equivalence highlights our method’s potential to achieve optimal performance by leveraging multiple non-conformity scores, even without a perfect classifier, approximating the performance of an ideal classifier through appropriate weight assignment.
6 Conclusion
In conclusion, our weighted aggregation approach addresses the challenge of constructing efficient and valid prediction sets in multi-class classification by learning optimal weights for combining different score functions using various data split procedures. This adaptability allows our method to adjust to the specific characteristics of the dataset, potentially outperforming methods that rely on a single score function. Our theoretical analysis establishes finite-sample guarantees for coverage, ensures the validity of the prediction sets, and derives oracle inequalities that quantify the expected set size of our approach relative to an optimal weight.
Experiments on the CIFAR-100 dataset demonstrate the superior performance of our weighted approach compared to single-score methods, consistently achieving the guaranteed coverage while maintaining the smallest prediction set sizes, which highlights the potential of our method to achieve optimal performance by leveraging the strengths of multiple non-conformity scores. Future research directions include exploring the application of our approach to other domains and investigating the integration of advanced machine learning techniques to further enhance the performance and efficiency of conformal prediction sets.
Appendix A Proof of Lemma 1
A.1 VC Dimension of the Subgraph Classes
Proposition 1.
Both of the following classes of functions
$\{(x,y)\mapsto\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq t\}:\bm{w}\in\mathbb{R}^{d},t\in\mathbb{R}\}$ (11)
and
$\{x\mapsto\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq t\}:\bm{w}\in\mathbb{R}^{d},t\in\mathbb{R}\},$ (12)
where $y\in[K]$ is fixed, have VC-dimension at most $d+1$.
Proof.
Both of the classes are the subgraph classes of vector space of functions with dimension $d+1$. By Proposition 4.20 of Wainwright (2019), these subgraph classes have dimension at most $d+1$. This is also a direct result of Example 4.21 in the book. ∎
A.2 Proof of the Lemma
This proof basically follows from Vapnik–Chervonenkis theory. We will use the theorems and lemmas in Wainwright (2019) as reference. The proof of Eq. (9) and Eq. (10) are almost identical. We will focus on Eq. (9). Let $\epsilon_{i},i\in\mathcal{I}$ be i.i.d. symmetric random variables takes value $-1$ or $1$, i.e., $\mathbb{P}(\epsilon_{i}=-1)=\mathbb{P}(\epsilon_{i}=1)=0.5.$ Then we define the Rademacher complexity for the function class in Eq. (11),
$\mathcal{R}(\mathcal{I}):=\mathbb{E}\left[\sup_{w,t}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\epsilon_{i}\mathbf{1}\{\langle w,s(X_{i},Y_{i})\rangle\geq t\}\right|\right].$
By Theorem 4.10 in the book,
$\mathbb{P}\left(\Omega\left(\mathcal{I},\mathcal{R}(\mathcal{I})+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (13)
By Proposition 1, the VC dimension of the function class in Eq. (11) is $d+1$. Then altogether with Lemma 4.14 and Proposition 4.18 in the book,
$\mathcal{R}(\mathcal{I})\leq 4\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}$ (14)
We combine Eq. (13) and Eq. (14) to obtain Eq. (9). The proof of Eq. (10) follows analogously, using the VC dimension result for the function class in Eq. (12). Let us define
$\Gamma_{y}(\mathcal{I},\eta)=\sup_{w\in\mathbb{R}^{d},t\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle w,s(X_{i},y)\rangle\geq t\}-\mathbb{E}[\mathbf{1}\{\langle w,s(X,y)\rangle\geq t\}]\right|\leq\eta.$
For $\mathcal{I}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$,
$\mathbb{P}\left(\Gamma_{y}\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$
It is clear that
$\bigcup_{y\in[K]}\Gamma_{y}(\mathcal{I},\eta)\subseteq\Gamma(\mathcal{I},K\eta).$
Taking the union bound on $y\in[K]$ implies Eq. (10).
A.3 Bounds for Coverage Probability
The following result requires the data for optimizing $\widehat{\bm{w}}$ are independent with the calibration set and the test set. The algorithm satisfying these conditions have the most reliable coverage probability.
Lemma 2.
Let $\widehat{C}_{1-\alpha}(x)$ be the output of Algorithm 2. Suppose
  • (i)
    $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable.
  • (ii)
    $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{1}\cup\mathcal{I}_{2}}$ and $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are independent.
Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X))\geq 1-\alpha.$
Proof.
$\widehat{\bm{w}}$ only depends on $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{1}\cup\mathcal{I}_{2}}$, so it is independent of $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$. The weighted score function $\{\langle\widehat{\bm{w}},s(X_{i},Y_{i})\rangle\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are also exchangeable. For any $(X,Y)$ in the test set, the rank of $s(X,Y)$ is smaller than $Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ with probability $\frac{\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil}{1+|\mathcal{I}_{3}|}\geq 1-\alpha$. ∎
Lemma 3.
Let $\widehat{C}_{1-\alpha}(x)$ be the output of Algorithm 2. Suppose
  • (i)
    $\Omega(\mathcal{I}_{3},\eta_{3})$ is satisfied.
  • (ii)
    $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{1}\cup\mathcal{I}_{2}}$ and $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\textrm{test}}}$ are independent.
Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X))\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}.$
Proof.
Let us write the threshold in Eq. (5) $Q_{1-\alpha}^{(2)}:=Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ to emphasize that this quantile depends on $\widehat{\bm{w}}$ and the samples in $\mathcal{I}_{3}$. By the procedure of the algorithm, we have
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}(X)\right)=\mathbb{P}\left(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\right).$
Assuming the event $\Omega(\mathcal{I}_{3},\eta_{3})$,
$\begin{split}Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})&=\sup\left\{t\in\mathbb{R}:\frac{1}{|\mathcal{I}_{3}|}\sum_{i\in\mathcal{I}_{3}}\mathbf{1}\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle\geq t\}\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil\right\}\\ &\geq\sup\left\{t\in\mathbb{R}:\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq t)\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}\right\}\\ &=Q_{1-\alpha^{\prime}}(\widehat{\bm{w}}),\end{split}$ (15)
where $1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}.$ By the definition of the quantile,
$\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3}))\geq\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha^{\prime}}(\widehat{\bm{w}}))=1-\alpha^{\prime}.$ (16)
The proof is complete. ∎
Lemma 4.
Let $\widehat{C}_{1-\alpha}(X;\widehat{\bm{w}})$ be the output of Algorithm 2. Suppose $\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}(x_{i})\}\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}-\eta_{\text{test}}.$
Proof.
Under the event $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$, for all $w\in\mathcal{W}$,
$\left|\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},Y_{i})\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}-\mathbb{E}_{(X,Y)}[\mathbf{1}\{\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}]\right|$
$\leq\sup_{\bm{w}\in\mathbb{R}^{d},t\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq t\}-\mathbb{E}_{(X,Y)}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq t\}]\right|\leq\eta_{\text{test}},$
where $(X,Y)$ is an i.i.d. copy of $(X_{i},Y_{i})$ in the test set. In particular, we let $\bm{w}=\widehat{\bm{w}}$ and $t=Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$, we have
$E[\mathbf{1}\{\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{3})\}\mid\widehat{\bm{w}}\ ]=\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\mid\ \widehat{\bm{w}})=1-\alpha.$
The remaining proof has similar argument as Eq. (15) and Eq. (16), and is omitted here. ∎
A.4 Bounds for Prediction Set Size
Lemma 5.
Suppose the samples in $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy $\Omega(\mathcal{I}_{1},\eta_{1})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ respectively, and suppose $Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{1})\leq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$, then for $X$ in the test set, the expected prediction set size satisfies
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha_{1}}(w^{*})\}\right]+\xi_{2},$
where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil+\eta_{1}$.
Proof.
Under the assumption $Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{1})\leq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$, for $w\in\mathcal{W}$, we have
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w,s(X_{i},y)\rangle\geq Q_{1-\alpha}(w,\mathcal{I}_{3})\}\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w,s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w,\mathcal{I}_{1})\}.$
$\widehat{\bm{w}}$ is obtained from
$\widehat{\bm{w}}\in\arg\min_{w\in\mathcal{W}}\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w,s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w,\mathcal{I}_{1})\}$
Therefore,
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{1})\}\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*},\mathcal{I}_{1})\}.$
Given the event $\Omega(\mathcal{I}_{1},\eta_{1})$, for $w\in\mathcal{W}$ and $t\in\mathbb{R}$,
$\mathbb{P}(\langle w,s(X,Y)\rangle\geq t)=\mathbb{E}[\mathbf{1}\{\langle w,s(X,Y)\rangle\geq t\}]\leq\frac{1}{|\mathcal{I}_{1}|}\sum_{i\in\mathcal{I}_{1}}\mathbf{1}\{\langle w,s(X,Y)\geq t\}+\eta_{1}.$
The lower bound of the LHS is also a lower bound of the RHS, so we have
$\begin{split}Q_{1-\alpha^{\prime}}(w^{*},\mathcal{I}_{1})&=\sup\left\{t\in\mathbb{R}:\frac{1}{|\mathcal{I}_{1}|}\sum_{i\in\mathcal{I}_{1}}\mathbf{1}\{\langle w,s(x_{i},y_{i})\rangle\geq t\}\geq\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil\right\}\\ &\geq\sup\left\{t\in\mathbb{R}:\mathbb{P}(\langle w^{*},s(X,Y)\rangle\geq t)\geq\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil+\eta_{1}\right\}\\ &=Q_{1-\alpha_{1}}(w^{*}),\end{split}$ (17)
where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil+\eta_{1}$ and $Q_{1-\alpha_{1}}(\bm{w}^{*})$ is the $(1-\alpha_{1})$-quantile of the population distribution $\langle w,s(X,Y)\rangle$. Now we can conclude that
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*},\mathcal{I}_{1})\}\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha_{1}}(w^{*})\}.$
On the event $\Gamma(\mathcal{I}_{2},\xi_{2})$,
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*})\}\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha_{1}}(w^{*})\}\right]+\xi_{2}.$
The proof is complete. ∎
Appendix B Proof of the Theorems
B.1 Proof of Theorem 1
Under the assumption of the theorem, the setting of VFCP satisfies the condition of Lemma 2. This proves the coverage probability in the theorem. For the expected prediction set size, under the event $\Gamma(\mathcal{I}_{2},\xi_{2})$, by Lemma 5, for $X$ in the test set,
$\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}\right]\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}+\xi_{2}.$
This implies the coverage probability of the theorem.
B.2 Proof of Theorem 2
Under the assumption of the theorem, the setting of VFCP satisfies the condition of Lemma 3. This proves the coverage probability in the theorem. For the expected prediction set size, since $\mathcal{I}_{1}=\mathcal{I}_{3}$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ is satisfied, by Lemma 5, for $X$ in the test set,
$\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}\right]\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}+\xi_{2}.$
This implies the coverage probability of the theorem.
B.3 Proof of Theorem 3
Since $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$, one can verify that the theorem is the direct result of Lemma 4 and Lemma 5.
B.4 Proof of Theorem 4
Similar as the setting of DLCP, the coverage probability is the direct result of Lemma 4. For the prediction set size, suppose $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for $X$ in the test set,
$\mathbb{E}[|\widehat{C}_{1-\alpha}^{\text{DLCP+}}(X)|]$ $\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}+\xi_{2}$
$\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*})\}\right]+2\xi_{2}.$

### v2

@@
-Weighted Aggregation of Conformity Scores for Classification
+Averaging of Conformity Scores for Classification
 Abstract
-Conformal prediction is a powerful framework for constructing prediction sets with valid coverage guarantees in multi-class classification.
-However, existing methods often rely on a single score function, which can limit their efficiency and informativeness.
-We propose a novel approach that combines multiple score functions to improve the performance of conformal predictors by identifying optimal weights that minimize prediction set size.
-Our theoretical analysis establishes a connection between the weighted score functions and subgraph classes of functions studied in Vapnik-Chervonenkis theory, providing a rigorous mathematical basis for understanding the effectiveness of the proposed method.
-Experiments demonstrate that our approach consistently outperforms single-score conformal predictors while maintaining valid coverage, offering a principled and data-driven way to enhance the efficiency and practicality of conformal prediction in classification tasks.
+Conformal prediction provides a robust framework for generating prediction sets with finite-sample coverage guarantees, independent of the underlying data distribution.
+However, existing methods typically rely on a single conformity score function, which can limit the efficiency and informativeness of the prediction sets.
+In this paper, we present a novel approach that enhances conformal prediction for multi-class classification by optimally averaging multiple conformity score functions.
+Our method involves assigning weights to different score functions and employing various data splitting strategies.
+Additionally, our approach bridges concepts from conformal prediction and model averaging, offering a more flexible and efficient tool for uncertainty quantification in classification tasks.
+We provide a comprehensive theoretical analysis grounded in Vapnik–Chervonenkis (VC) theory, establishing finite-sample coverage guarantees and demonstrating the efficiency of our method.
+Empirical evaluations on benchmark datasets show that our weighted averaging approach consistently outperforms single-score methods by producing smaller prediction sets without sacrificing coverage.
 1 Introduction
-In the era of data-driven decision-making, the ability to quantify uncertainty in machine learning predictions is becoming increasingly important, particularly in high-stakes applications such as medical criminal justice (Berk et al., 2021) and diagnostics (Angelopoulos et al., 2024).
-While most machine learning methods focus on providing point predictions, supplementing these predictions with valid uncertainty estimates in the form of prediction sets or intervals, is crucial for responsible decision-making.
-Conformal prediction (Vovk et al., 2005; Manokhin, 2022) is a powerful framework that provides valid prediction sets with finite-sample coverage guarantees, regardless of the true underlying distribution (Angelopoulos et al., 2023).
-The key idea is to construct a prediction set for a new test point based on training data, such that the probability of the true label being contained in the prediction set is at least $1-\alpha$.
-This coverage guarantee holds irrespective of the point prediction algorithm used, making conformal prediction a versatile tool for quantifying uncertainty in machine learning predictions.
-Central to conformal prediction is the concept of score functions, which quantify the conformity of a data point to the underlying distribution.
-In the split conformal prediction framework (Papadopoulos et al., 2002; Lei & Wasserman, 2014; Vovk et al., 2018), the training data is divided into a proper training set and a calibration set.
-The point prediction algorithm is trained on the proper training set, while the score function is evaluated on the calibration set.
-The conformal prediction set is then constructed by including all points whose score falls below a certain quantile of the calibration scores, with the quantile determined by the desired coverage level $1-\alpha$.
-The choice of score function plays a crucial role in determining the efficiency of the resulting prediction sets, with well-chosen score functions leading to more informative and precise predictions.
-The flexibility in choosing score functions allows conformal prediction to adapt to the specific characteristics of the data and the underlying distribution.
-This is an active area of research in conformal prediction, with ongoing efforts to develop score functions that yield the most informative and efficient prediction sets for a given problem setting, including regression (Papadopoulos et al., 2008, 2011; Romano et al., 2019; Kivaranovic et al., 2020; Guan, 2023; Colombo, 2023, 2024) and classification (Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2021; Huang et al., 2024; Luo & Zhou, 2024).
-Despite the dependence of prediction set efficiency on the choice of score function and the performance of the underlying point prediction algorithm, the coverage guarantee of conformal prediction remains intact, making it a robust tool for providing reliable uncertainty estimates in machine learning predictions.
-In this work, we propose a novel approach to optimize a linear combination of conformity scores using efficiency as the objective function.
-Our goal is to search for optimal weights for such linear combinations, enabling us to fully utilize the advantages of existing score functions.
-While our approach shares similarities with Izbicki et al.
-(2019) and Yang & Kuchibhotla (2024), which discuss the construction of the smallest prediction set using kernel density estimators and families of machine learning algorithms, respectively, our method distinguishes itself in three key aspects:
+Conformal prediction Vovk et al.
+(2005); Manokhin (2022) is a robust framework that generates prediction sets with finite-sample coverage guarantees, irrespective of the underlying data distribution Angelopoulos et al.
+(2023).
+The fundamental principle of conformal prediction is to construct a prediction set for a new test instance based on the training data, ensuring that the true label is included with a probability of at least $1-\alpha$.
+This coverage assurance holds regardless of the specific point prediction algorithm employed, making conformal prediction a versatile tool for uncertainty quantification in machine learning.
+In the split conformal prediction framework Papadopoulos et al.
+(2002); Lei and Wasserman (2014); Vovk et al.
+(2018), the training data is partitioned into a training set and a calibration set.
+The predictive model is trained on the training subset, while the score functions are evaluated on the calibration subset.
+The conformal prediction set then comprises all labels whose scores fall below a specific quantile of the calibration scores, with the quantile determined by the coverage level.
+The choice of score function is critical in determining the efficiency of the resulting prediction sets, especially for multi-class classification.
+Well-chosen score functions can lead to more informative and precise predictions.
+This flexibility allows conformal prediction to adapt to the specific characteristics of the data and distribution.
+Consequently, developing score functions that optimize informativeness and efficiency for various problem settings, including regression Papadopoulos et al.
+(2008, 2011); Romano et al.
+(2019); Kivaranovic et al.
+(2020); Guan (2023); Colombo (2023, 2024) and multi-class classification Sadinle et al.
+(2019); Romano et al.
+(2020); Angelopoulos et al.
+(2021); Huang et al.
+(2024); Luo and Zhou (2024), remains an active area of research.
+This work focuses on enhancing the efficiency of prediction sets for classification tasks.
+While the underlying idea can extend to regression, this paper will focus on the tasks of multi-class classification.
+Our approach assumes the availability of multiple score functions for the same classification task, each differing due to variations in the classification algorithm or the definition of the score.
+We propose assigning optimal weights to aggregate these score functions.
+Using a validation set, we determine a threshold to achieve the desired coverage and identify the weight combination that minimizes the prediction set size.
+The final prediction is then based on this weighted score function.
+Our aim is to find the optimal weights for linear combinations of score functions, thereby fully leveraging the strengths of existing score functions.
+While our approach shares similarities with Yang and Kuchibhotla (2024), it stands out in three key aspects:
   1. 1.
-    First, our approach emphasizes finding an optimal combination of score functions, which is more practical in the current landscape of numerous high-quality and diverse score functions, rather than focusing on training multiple algorithms.
-By leveraging the strengths of existing score functions, we aim to improve the efficiency of conformal prediction while maintaining its coverage guarantees.
+    Weighted Averaging of Score Functions: Instead of selecting the single best-performing score function, our approach combines multiple score functions through optimal weighting.
+This averaging can yield more efficient prediction sets than any individual score function while maintaining the desired coverage guarantees.
   2. 2.
-    Second, we investigate and categorize major data split strategies for determining optimal weights for combining score functions.
-These strategies include Validity First Conformal Prediction (VFCP) and Efficiency First Conformal Prediction (EFCP) (Yang & Kuchibhotla, 2024), Data Leakage Conformal Prediction (DLCP), and its variant that utilizes all available data (DLCP+).
-We establish the coverage property for VFCP and quantify the coverage gap for the other three strategies, along with their corresponding efficiency improvements.
+    Novel Data Splitting Strategies: We explore and categorize several data splitting methods to determine the optimal weights for combining score functions.
+In addition to Validity First Conformal Prediction (VFCP) and Efficiency First Conformal Prediction (EFCP) discussed in Yang and Kuchibhotla (2024), we introduce Data Leakage Conformal Prediction (DLCP) and its variant DLCP+, which utilize all available data to enhance weight determination.
   3. 3.
-    Third, our method distinguishes itself from selection methods that focus on identifying the single best-performing score function.
-By aggregating multiple score functions through an optimally chosen weight, our approach can achieve performance levels that surpass those of any individual score function.
-This ability to harness the collective strengths of multiple score functions sets our method apart and enables it to provide more efficient and informative prediction sets.
-Our weighted aggregation approach addresses the challenge of constructing efficient and valid prediction sets in multi-class classification by learning optimal weights for combining different score functions using various data split procedure.
-This allows for adaptation to the specific characteristics of the dataset and potentially outperforms methods relying on a single score function.
-We provide a theoretical analysis of our approach, establishing finite-sample guarantees for coverage and width and deriving oracle inequalities to quantify its performance relative to an “optimal” aggregation.
-Through extensive experiments, we demonstrate the superior performance of our weighted aggregation approach compared to single-score methods, highlighting its potential for improving the efficiency of conformal prediction in classification tasks.
-The rest of the paper is organized as follows.
-In Section 2, we present our weighted aggregation approach and the various data split strategies.
-Section 3 contains our theoretical analysis, establishing coverage guarantees and deriving the expected prediction set sizes.
-In Section 4, we discuss our experimental results, demonstrating the effectiveness of our approach.
-Related works are presented in Section 5.
-Finally, we conclude the paper in Section 6 and discuss future research directions.
-2 Weighted Aggregation of Score Functions
-2.1 Basic Conformal Prediction for Classification
-We start by assuming that a classification algorithm provides $\widehat{p}_{y}(x)$, which approximates $P(Y=y|X=x)$.
+    Theoretical Foundations Using VC Theory: We provide a theoretical analysis of our method that leverages Vapnik–Chervonenkis theory to establish coverage guarantees and expected prediction set sizes.
+This solid mathematical foundation underscores the validity and efficiency of our approach across different data splitting strategies.
+Beyond its novel contributions to the conformal prediction literature, our method is closely related to model averaging Claeskens and Hjort (2008), a well-established technique in machine learning.
+Unlike traditional model averaging, which assigns weights to different models to improve prediction accuracy, our method assigns weights to score functions.
+This distinction requires the development of specific data splitting techniques to ensure the desired coverage guarantees.
+Consequently, our work can be viewed as an innovative adaptation of model averaging principles to the conformal prediction framework.
+The remainder of the paper is organized as follows.
+In Section 3, we detail our weighted averaging approach and the various data splitting strategies employed.
+Section 3 presents the theoretical analysis, establishing coverage guarantees and expected prediction set sizes.
+In Section 4, we demonstrate the effectiveness of our method through experiments.
+Related works are discussed in Section 5.
+We conclude in Section 6 and outline future research directions.
+2 Methodology
+2.1 Conformal Prediction for Classification
+We start by assuming that a $K$-class classification algorithm provides $\widehat{p}_{y}(x)$, which approximates $P(Y=y|X=x)$ for $y\in[K]$.
 While our method and theoretical analysis do not depend on the accuracy of this approximation, it is beneficial to assume that higher values of $\widehat{p}_{y}(x)$ indicate a greater likelihood of sample $x$ having label $y$.
-We consider this training procedure to be performed on a separate dataset, ensuring that $\widehat{p}(y|x)$ is independent of the dataset used in this paper.
+We consider this training procedure to be performed on a separate dataset, ensuring that $\widehat{p}_{y}(x)$ is independent of the dataset used in this paper.
 Let us first consider a single non-conformity score function $s(x,y)$.
@@
 In this context, conformal prediction for classification problems can be described by the following algorithm.
-To summarize, we first find a threshold in a set of labeled data, so that $s(x_{i},y_{i})\geq Q_{1-\alpha}$ holds for at least $1-\alpha$ proportion in the set $\mathcal{I}_{\textrm{train}}$.
+To summarize, we first find a threshold in a set of labeled data, so that $s(x_{i},y_{i})\geq Q_{1-\alpha}$ holds for at least $1-\alpha$ proportion in the set $\mathcal{I}_{2}$.
 Then we use this threshold to define the prediction set for $x_{i}$ in the test set, which is the upper level set of the function $y\mapsto s(x_{i},y)$.
+This algorithm constructs prediction sets $\widehat{C}_{1-\alpha}(x_{i})$ for each $i\in\mathcal{I}_{\textrm{test}}$, based on the non-conformity scores and a threshold determined by the desired coverage probability $1-\alpha$.
+The algorithm splits the training data into two subsets: $\mathcal{I}_{1}$ for calculating the non-conformity scores, and $\mathcal{I}_{2}$ for calibration.
+The threshold $Q_{1-\alpha}$ is chosen to ensure the desired coverage probability under the exchangeability assumption.
 Algorithm 1 Split Conformal Prediction
@@
 6:$\widehat{C}_{1-\alpha}(x_{i},Q_{1-\alpha})\leftarrow\{y\in[K]:s(x_{i},y)\geq Q_{1-\alpha}\}$ for $i\in\mathcal{I}_{\textrm{test}}$.
-This algorithm constructs prediction sets $\widehat{C}_{1-\alpha}(x_{i})$ for each $i\in\mathcal{I}_{\textrm{test}}$, based on the non-conformity scores and a threshold determined by the desired coverage probability $1-\alpha$.
-The algorithm splits the training data into two subsets: $\mathcal{I}_{1}$ for calculating the non-conformity scores, and $\mathcal{I}_{2}$ for calibration.
-The threshold $Q_{1-\alpha}$ is chosen to ensure the desired coverage probability.
-A key property of this method is that, under the assumption of exchangeability of the data points (a weaker assumption than i.i.d.), it guarantees that the resulting prediction sets will contain the true label with probability at least $1-\alpha$.
-This property holds regardless of the accuracy of the underlying classification algorithm, making conformal prediction a powerful tool for uncertainty quantification.
-2.2 Example of Score Functions for Classifications
-In the context of conformal prediction, the choice of score function $s(x,y)$ is crucial as it directly impacts the resulting prediction sets.
-These score functions serve as non-conformity measures, quantifying how well a label fits a given input.
-Let $K$ denote the number of classes in our classification problem.
-Various score functions have been proposed and their effectiveness demonstrated through extensive experimentation on diverse datasets. Three typical choices of score functions are:
+$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\text{train}}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{3}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{2}}$ $Q_{1-\alpha}^{(1)}(\bm{w})$ by Eq.
+(2) $\big{(}\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$ by Eq.
+(3) $\widehat{\bm{w}}$ by Eq.
+(4) $Q_{1-\alpha}^{(2)}$ by Eq.
+(5) $\{x_{i}\}_{i\in\mathcal{I}_{\text{test}}}$ $\left(\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\right)_{i\in\mathcal{I}_{\text{test}}}$ by Eq.
+(6)
+Figure 1: This example illustrates a framework for data splitting into $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$, and $\mathcal{I}_{\text{test}}$.
+Algorithm 2 presents the complete procedure.
+Briefly, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are used in Steps 1-2 to select the optimal weight $\widehat{\mathbf{w}}$, while $\mathcal{I}_{3}$ is used in Step 3 as the calibration set for $\mathcal{I}_{\text{test}}$ predictions.
+We present four options: VFCP, EFCP, DLCP, and DLCP+.
+Their coverage and size properties are discussed theoretically in Section 3 and empirically in Section 4.
+2.2 Various Score Functions for Classifications
+In conformal prediction, the choice of score function $s(x,y)$ critically influences the prediction sets.
+These functions act as non-conformity measures that quantify how appropriately a label matches an input.
+Score functions are determined by two primary factors: (1) The predictive model’s quality in estimating $\widehat{p}_{y}(x)\approx\mathbb{P}(Y=y|X=x)$, and (2) The specific methodology for converting these probability estimates into scores.
+We will consider approaches:
   1. 1.
-    Threshold (THR) (Sadinle et al., 2019): $s_{1}(x,y)=\widehat{p}_{y}(x)$.
+    Threshold (THR) Sadinle et al. (2019): $s_{\text{THR}}(x,y)=\widehat{p}_{y}(x)$
   2. 2.
-    Adaptive Prediction Sets (APS) (Romano et al., 2020): $s_{2}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\}$.
+    Adaptive Prediction Sets (APS) Romano et al. (2020): $s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\}$
   3. 3.
-    Rank-based (RANK) (Luo & Zhou, 2024): $s_{3}(x,y)=|\{y^{\prime}\in[K]:\widehat{p}_{y^{\prime}}(x)<\widehat{p}_{y}(x)\}|$.
-We can easily plug the APS and RANK score functions into Algorithm 1 to create prediction sets.
-While the original publications introducing these score functions presented additional complexities to guarantee precise $1-\alpha$ expected coverage, these adjustments do not significantly impact the presentation of our proposed approach and are therefore omitted for clarity.
-The reasoning behind these scores will not affect how we combine them or our theoretical analysis.
-In the following sections, we will demonstrate how these score functions can be effectively utilized in our proposed approach without the need for additional complexity to ensure exact coverage.
-A detailed introduction and discussion of these score functions, along with their adjustments, will be provided in Section 4.
-2.3 Aggregation of Score Functions
-In Yang & Kuchibhotla (2024), the authors propose aggregating score functions for conformal prediction.
+    Rank-based (RANK) Luo and Zhou (2024): $s_{\text{RANK}}(x,y)=|\{y^{\prime}\in[K]:\widehat{p}_{y^{\prime}}(x)<\widehat{p}_{y}(x)\}|$
+These score functions integrate directly with Algorithm 1.
+While the original formulations include additional refinements to achieve exact $1-\alpha$ coverage guarantees, these refinements do not materially affect our combination methodology and are omitted for simplicity.
+Our theoretical analysis remains invariant to specific score function design choices.
+Other notable approaches like RAPS Angelopoulos et al.
+(2021) and SAPS Huang et al.
+(2024) demonstrate alternative methodologies for deriving scores from $\widehat{p}_{y}(x)$.
+2.3 Averaging Score Functions
+In Yang and Kuchibhotla (2024), the authors propose their method with multiple score functions.
 Their approach selects the score function that yields the smallest average prediction set size among the available score functions.
 They also introduce two data splitting methods: Efficiency First Conformal Prediction (EFCP) and Validity First Conformal Prediction (VFCP).
-Given a vector of score functions $s(x,y)=(s_{1}(x,y),\dots,s_{d}(x,y))^{\top}$, instead of simply choosing the score that provides the smallest prediction set on a calibration set, we propose assigning weights to the scores and defining a new weighted score function:
+Given a vector of score functions $s(x,y)=(s_{1}(x,y),\dots,s_{d}(x,y))^{\top}$, instead of simply choosing the score that provides the smallest prediction set on a calibration set, we propose assigning weights $\bm{w}\in\mathcal{W}\subseteq\mathbb{R}^{d}$ to the non-confomity scores and defining the weighted score function:
 $\langle\bm{w},s(x,y)\rangle=\sum_{j=1}^{d}\bm{w}_{j}s_{j}(x,y).$ (1)
@@
 This approach allows for more flexibility in combining the strengths of various score functions, potentially leading to improved prediction set efficiency.
-2.4 Algorithm for Determining the Weight and the Threshold
-In this section, we present a detailed procedure for determining the optimal weight vector $\bm{w}$ in the context of the weighted score approach introduced in Eq.
-(1).
-This procedure extends Algorithm 1 by incorporating an additional step to minimize the expected size of the prediction set.
-The following steps outline the process:
-  1. 1.
-    Extract $\mathcal{I}_{1},\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ from the dataset.
-In the meta version of the algorithm, these sets can be arbitrary, provided they are sufficiently large.
-However, for clarity and ease of understanding for first-time readers, it is helpful to assume that $\mathcal{I}_{1}=\mathcal{I}_{2}$ and that $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are partitions of $\mathcal{I}_{\textrm{train}}$.
-  2. 2.
-    Define a finite subset $\mathcal{W}\subseteq\mathbb{R}^{d}$ for candidates of $\bm{w}$.
-Theoretically, $\mathcal{W}$ can be an arbitrary subset of $\mathbb{R}^{d}$.
-Given that the scores are all non-conformity scores, it is sometimes desirable for the weights to be non-negative.
-In this case, letting $\mathcal{W}$ be the grid of the simplex $\Delta(d)\subseteq\mathbb{R}^{d}$ is a typical choice.
-The compactness of $\Delta(d)$ can make the algorithm computationally feasible.
-  3. 3.
-    Estimate the threshold for every $\bm{w}\in\mathcal{W}$ on $\mathcal{I}_{1}$. In the set $\mathcal{I}_{1}$, calculate the threshold for every $\bm{w}\in\mathcal{W}$:
-    $Q_{1-\alpha}^{(1)}(\bm{w})=\lceil(1+|\mathcal{I}_{1}|)(1-\alpha)\rceil\text{-th smallest value of }\{\langle\bm{w},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{1}\}.$ (2)
-    This corresponds to Step 5 in Algorithm 1.
-  4. 4.
-    Determine the prediction set for data in $\mathcal{I}_{2}$.
-For a given $\bm{w}$ and threshold $Q_{1-\alpha}^{(1)}(\bm{w})$, define the prediction set for each sample $x_{i},i\in\mathcal{I}_{2}$ as:
-    $\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))=\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(1)}(\bm{w})\}\quad\text{for}\quad i\in\mathcal{I}_{2}.$ (3)
-    This corresponds to Step 6 in Algorithm 1.
-However, we are not done yet, since these are the prediction sets for all $\bm{w}\in\mathcal{W}$, and our goal is to minimize the prediction set size.
-This intuitively leads to the next step.
-  5. 5.
-    Find the optimal weight vector $\widehat{\bm{w}}$ that minimizes the average prediction set size:
-    $\widehat{\bm{w}}\in\arg\min_{\bm{w}\in\mathcal{W}}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))|.$ (4)
-  6. 6.
-    If $\mathcal{I}_{3}\neq\mathcal{I}_{1}$, then $\mathcal{I}_{3}$ is treated as the calibration set in Step 5 in Algorithm 1, and we obtain:
-    $Q_{1-\alpha}^{(2)}=\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil\text{-th smallest value of }\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{3}\}.$ (5)
-    If $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are disjoint, $\widehat{\bm{w}}$ and $Q_{1-\alpha}^{(2)}$ are independent.
-If $\mathcal{I}_{1}=\mathcal{I}_{3}$, then this quantile has been computed in Eq.
-(2), i.e., $Q_{1-\alpha}^{(2)}=Q_{1-\alpha}^{(1)}(\widehat{\bm{w}})$.
-  7. 7.
-    The final output is the confidence set:
-    $\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})=\{y\in[K]:\langle\widehat{\bm{w}},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(2)}\}\quad\text{for}\quad i\in\mathcal{I}_{\textrm{test}}.$ (6)
-The following two key observations of this procedure:
-  • (a)
-    Once the optimal weight vector $\widehat{\bm{w}}$ has been determined, the subsequent steps of the calibration procedure, as described in Eq.
-(5), and the construction of the prediction set, as outlined in Eq.
-(6), are identical to Step 5 and Step 6 of Algorithm 1.
-The additional steps preceding these are designed to identify an appropriate weight vector $\bm{w}$ that optimizes the performance of the algorithm.
-  • (b)
-    The operation in Eq.
-(2) and Eq.
-(5) are the steps for finding quantiles, and Eq.
-(3) and Eq.
-(6) are the steps for defining prediction sets.
-For finding the quantiles, we need the labels of the samples, while for defining the prediction set, only the feature $x$ of the sample is required.
-In other words, it is possible to include the test set in Eq.
-(3) and Eq.
-(6).
-In the following algorithm, we present a formalized version of the procedure outlined above.
 Algorithm 2 Optimal Weight Selection and Prediction Set Construction
-1:$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$: Labeled training data; $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$: Unlabeled test data; $\alpha\in(0,1)$: Pre-defined significance level; $\mathcal{W}\subseteq\mathbb{R}^{d}$: Set of possible weight vectors, where $d$ is the number of score functions.
-2:$\big{(}C_{1-\alpha}(X_{i})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$: Sequence of prediction sets for test data with anticipated $1-\alpha$ coverage
+1:$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$: Labeled training data; $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$: Unlabeled test data; $\alpha\in(0,1)$: Pre-defined significance level; $s_{i}(\cdot,\cdot),i=1,\dots,d$: score functions.
+$\mathcal{W}\subseteq\mathbb{R}^{d}$: Set of possible weight vectors.
+2:$\big{(}C_{1-\alpha}(X_{i})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$: Sequence of prediction sets for test data with anticipated $1-\alpha$ coverage.
 3:Decide $\mathcal{I}_{1}\subseteq\mathcal{I}_{\textrm{train}}$, $\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$, and $\mathcal{I}_{3}\subseteq\mathcal{I}_{\textrm{train}}$ for Step 1, 2, and 3, respectively
@@
 9:for $\bm{w}\in\mathcal{W}$ do
-10:     $\big{(}\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}\leftarrow$ Evaluation($\bm{w},\mathcal{I}_{2},Q_{1-\alpha}^{(1)}(\bm{w})$).
-11:     Compute average prediction set size $S{(\bm{w})}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w})\big{)}|$.
-12:end for
-13:$\widehat{\bm{w}}\leftarrow\arg\min_{\bm{w}\in\mathcal{W}}S{(\bm{w})}$.
-14:Step 3: Test-time Prediction Set Construction
-15:$Q_{1-\alpha}^{(2)}\leftarrow$ Calibration($\widehat{\bm{w}},\mathcal{I}_{3},\alpha$).
-16:$\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}\leftarrow$ Evaluation($\widehat{\bm{w}},\mathcal{I}_{\textrm{test}},\mathcal{I}_{3}$).
-17:return $\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$
+10:     $\big{(}\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$
+11:             $\leftarrow$ Evaluation($\bm{w},\mathcal{I}_{2},Q_{1-\alpha}^{(1)}(\bm{w})$).
+12:     Compute average prediction set size $S{(\bm{w})}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w})\big{)}|$.
+13:end for
+14:$\widehat{\bm{w}}\leftarrow\arg\min_{\bm{w}\in\mathcal{W}}S{(\bm{w})}$.
+15:Step 3: Test-time Prediction Set
+16:$Q_{1-\alpha}^{(2)}\leftarrow$ Calibration($\widehat{\bm{w}},\mathcal{I}_{3},\alpha$).
+17:$\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$leftarrow Evaluation($\widehat{\bm{w}},\mathcal{I}_{\textrm{test}},\mathcal{I}_{3}$).
+18:return $\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$
+2.4 The Optimal Weight and the Threshold
+In this section, we present a detailed procedure for determining the optimal weight vector $\bm{w}$ in the context of the weighted score approach introduced in Eq.
+(1).
+This procedure extends Algorithm 1 by incorporating an additional step to minimize the expected size of the prediction set.
+The following steps outline the process:
+  1. 1.
+    Extract $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ from dataset.
+Assume $\mathcal{I}_{1}=\mathcal{I}_{2}$ and $\mathcal{I}_{1},\mathcal{I}_{3}$ partition $\mathcal{I}_{\textrm{train}}$ for clarity.
+  2. 2.
+    Define $\mathcal{W}\subseteq\mathbb{R}^{d}$ for $\bm{w}$ candidates. For non-negative weights, use grid of simplex $\Delta(d)\subseteq\mathbb{R}^{d}$.
+  3. 3.
+    Estimate the threshold for every $\bm{w}\in\mathcal{W}$ on $\mathcal{I}_{1}$. In the set $\mathcal{I}_{1}$, calculate the threshold for every $\bm{w}\in\mathcal{W}$:
+    $Q_{1-\alpha}^{(1)}(\bm{w})=\ \lceil(1+|\mathcal{I}_{1}|)(1-\alpha)\rceil\text{-th smallest}\text{value of}\quad\{\langle\bm{w},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{1}\}.$ (2)
+    This corresponds to Step 5 in Algorithm 1.
+  4. 4.
+    Determine the prediction set for data in $\mathcal{I}_{2}$.
+For a given $\bm{w}$ and threshold $Q_{1-\alpha}^{(1)}(\bm{w})$, define the prediction set for each sample $x_{i},i\in\mathcal{I}_{2}$ as:
+    $\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))=\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(1)}(\bm{w})\}.$ (3)
+    This corresponds to Step 6 in Algorithm 1.
+However, we are not done yet, since these are the prediction sets for all $\bm{w}\in\mathcal{W}$, and our goal is to minimize the prediction set size.
+This intuitively leads to the next step.
+  5. 5.
+    The optimal weight vector $\widehat{\bm{w}}$ is obtained by minimizing the empirical prediction set size:
+    $\widehat{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\mathcal{W}}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))|,$ (4)
+    where $\mathcal{W}$ is a discrete grid over the $(d-1)$-dimensional probability simplex $\Delta^{d-1}$.
+We systematically explore candidate weights via grid search with step size $\varepsilon=0.01$, as detailed in Appendix A.
+  6. 6.
+    If $\mathcal{I}_{3}\neq\mathcal{I}_{1}$, then $\mathcal{I}_{3}$ is treated as the calibration set in Step 5 in Algorithm 1, and we obtain:
+[…]
+    value of $\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{3}\}.$
+    If $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are disjoint, $\widehat{\bm{w}}$ and $Q_{1-\alpha}^{(2)}$ are independent.
+If $\mathcal{I}_{1}=\mathcal{I}_{3}$, then this quantile has been computed in Eq.
+(2), i.e., $Q_{1-\alpha}^{(2)}=Q_{1-\alpha}^{(1)}(\widehat{\bm{w}})$.
+  7. 7.
+    The final output is the confidence set for samples $i\in\mathcal{I}_{\textrm{test}}$ in the test set:
+    $\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})=\{y\in[K]:\langle\widehat{\bm{w}},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(2)}\}.$ (6)
+This procedure comprises two primary steps: Threshold Calibration: Determining appropriate thresholds using Equations Eq.
+(2) and Eq.
+(5) based on the calibration subsets $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$.
+Prediction Set Evaluation: Constructing the prediction sets for $\mathcal{I}_{2}$ and $\mathcal{I}_{\text{test}}$ as outlined in Equations Eq.
+(3) and Eq.
+(6).
+These steps are encapsulated in Function 1 and Function 2, respectively.
+The entire procedure of the proposed method is summarized in Algorithm 2.
 Function 1 Calibration
@@
 2:     for $i\in\mathcal{I}_{\textrm{calib}}$ do
-3:         Compute score $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
+3:         Compute $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
 4:     end for
@@
 3:         for $y\in[K]$ do
-4:              Compute non-conformity score $\langle\bm{w},s(x_{i},y)\rangle$ using $\bm{w}$ and $(x_{i},y)$
+4:              Compute weighted score $\langle\bm{w},s(x_{i},y)\rangle$
 5:         end for
@@
 9:end function
-For $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$, we present an overview of the data split framework in Figure 1.
-For what follows, we introduce the following four choices.
+2.5 Data Splitting
+In Algorithm 2, the method for splitting the data into $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ has not been specified.
+To explore potential data splitting approaches, we first highlight the following two key observations.
+Firstly, After determining $\widehat{\bm{w}}$, the calibration procedure Eq.
+(5) and prediction set construction Eq.
+(6) mirror Step 5 and Step 6 of Algorithm 1.
+Prior steps aim to identify a weight vector $\bm{w}$ optimizing algorithm performance.
+Secondly, Eq.
+(2) and Eq.
+(5) find quantiles, requiring sample labels.
+Equations Eq.
+(3) and Eq.
+(6) define prediction sets, needing only feature $x$.
+Thus, the test set can be included in Eq.
+(3) and Eq.
+(6).
+Based on the observations above, we introduce the following four possible ways of data splitting.
   • (a)
-    For Validity First Conformal Prediction (VFCP) (Yang & Kuchibhotla, 2024), $\mathcal{I}_{1}=\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}\setminus\mathcal{I}_{1}$.
-This method divides $\mathcal{I}_{\textrm{train}}$ into two partitions, using less data to determine $\widehat{\bm{w}}$.
-However, the marginal coverage probability is guaranteed to be at least $1-\alpha$ under the assumption of exchangeability of samples in $\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$.
-The name “validity first” comes from this fact.
+    Validity First Conformal Prediction (VFCP) Yang and Kuchibhotla (2024): $\mathcal{I}_{1}=\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}\setminus\mathcal{I}_{1}$.
+Divides $\mathcal{I}_{\textrm{train}}$ into two partitions.
+Marginal coverage probability is guaranteed to be at least $1-\alpha$ under exchangeability of samples in $\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$.
   • (b)
-    For Efficiency First Conformal Prediction (EFCP) (Yang & Kuchibhotla, 2024), $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$.
-This method uses all training data to determine $\widehat{\bm{w}}$, resulting in a more accurate estimation of the optimal $\bm{w}$.
-However, stronger assumptions are required for the coverage guarantee.
-The name “efficiency first” comes from this fact.
+    Efficiency First Conformal Prediction (EFCP) Yang and Kuchibhotla (2024): $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$.
+Uses all training data to determine $\widehat{\bm{w}}$, resulting in more accurate estimation of optimal $\bm{w}$.
+This method equires stronger assumptions for coverage guarantee.
   • (c)
-    For Data Leakage Conformal Prediction (DLCP), $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$.
-Since $\mathcal{I}_{2}$ can be unlabeled data, we can minimize the prediction set on $\mathcal{I}_{\textrm{test}}$ in the step of finding $\widehat{\bm{w}}$ in Eq.
-(4) by letting $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$.
-We call this a “data leakage” method because we use the test data in the training procedure of $\widehat{\bm{w}}$.
+    Data Leakage Conformal Prediction (DLCP): $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$.
+Minimizes prediction set on $\mathcal{I}_{\textrm{test}}$ in finding $\widehat{\bm{w}}$ in Eq.
+(4).
+Called ”data leakage” as it uses test data in training procedure of $\widehat{\bm{w}}$.
   • (d)
-    For another version of Data Leakage Conformal Prediction (DLCP+), $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$.
-If data leakage is allowed, then using the same idea as EFCP, we can use all available data in each step.
-Therefore, we can include $\mathcal{I}_{\textrm{test}}$ in $\mathcal{I}_{2}$ to maximize the size of samples in the step of finding $\widehat{\bm{w}}$ in Eq.
+    Data Leakage Conformal Prediction+ (DLCP+): $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$.
+Uses all available data in each step, including $\mathcal{I}_{\textrm{test}}$ in $\mathcal{I}_{2}$ to maximize sample size for finding $\widehat{\bm{w}}$ in Eq.
 (4).
-$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{3}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{2}}$ $Q_{1-\alpha}^{(1)}(\bm{w})$ by Eq.
-(2) $\big{(}\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$ by Eq.
-(3) $\widehat{\bm{w}}$ by Eq.
-(4) $Q_{1-\alpha}^{(2)}$ by Eq.
-(5) $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$ $\left(\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\right)_{i\in\mathcal{I}_{\textrm{test}}}$ by Eq.
-(6)
-Figure 1: This example shows an overall framework of data splitting for $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$.
-The complete procedure is presented in Algorithm 2.
-Shortly speaking, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are used in Step 1 and 2 to select optimal weight $\widehat{\bm{w}}$, $\mathcal{I}_{3}$ is used in Step 3 as calibration set to construct prediction sets for $\mathcal{I}_{\textrm{test}}$.
-Under this framework, we illustrate four options, VFCP, EFCP, DLCP, and DLCP+, and discuss their coverage and size property in Section 3 as well as empirical results in Section 4.
+There are various possible methods for data splitting.
+In the following sections, we focus on these four specific approaches, examining their theoretical properties and evaluating their performance through experiments.
 3 Theoretical Analysis
-The theoretical analysis of our method are based on the subgraph classes in Vapnik–Chervonenkis (VC) theory.
-This theoretical details and the proofs of lemma and theorems will be left in Section A and B in the Appendix.
-In this section, we will present the key result for our proposed algorithm.
-The output of the algorithm analyzed in two aspect:
-  1. 1.
-    Can the prediction set achieve $1-\alpha$ coverage?
-  2. 2.
-    Can the prediction set achieve the minimal size? The minimal size will be defined as below.
-$\bm{w}^{*}$ is defined as the population level of $\widehat{\bm{w}}.$
-$\bm{w}^{*}=\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}[|\widehat{C}_{1-\alpha}(X,\bm{w},q)|]\quad{s.t.}\quad\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha.$
+3.1 Overview of the Results
+We investigate the theoretical properties of the proposed methods in terms of validity and efficiency.
+Validity: We assess whether the output prediction set achieves the desired coverage rate of $1-\alpha$.
+For the VFCP method, the coverage rate is guaranteed under the exchangeability assumption.
+However, for the other three methods, establishing validity is more complex due to the selection bias introduced by $\widehat{\bm{w}}$.
+Consequently, our aim is to demonstrate that the optimization over $\bm{w}\in\mathcal{W}$ has small impact on validity.
+Efficiency: We first observe that the prediction set attains the smallest possible expected size only if the true conditional probabilities $p(y|x)$ are known.
+Given the predefined score functions, the optimal weight vector in the population sense is defined as
+$\bm{w}^{*}=$ $\ \arg\min_{\bm{w}\in\mathcal{W}}\min_{q\in\mathbb{R}}\mathbb{E}\left[|\widehat{C}_{1-\alpha}(X,\bm{w},q)|\right]$
+s.t. $\quad\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q)\right)\geq 1-\alpha.$
+$\bm{w}^{*}$ is the weight vector that minimizes the expected prediction set size while ensuring that, together with an appropriate threshold $q$, the prediction set maintains a desired coverage rate of $1-\alpha$.
 Equivalently, we can define
-$Q_{1-\alpha}(\bm{w})=\sup\{q\in\mathbb{R}:P(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha\},$
+$Q_{1-\alpha}(\bm{w}):=\sup\{q\in\mathbb{R}:\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha\},$
 and using this definition, we define
-$\bm{w}^{*}=\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}[|\widehat{C}_{1-\alpha}(X,\bm{w},Q_{1-\alpha}(\bm{w})|]$
-Let us define $q^{*}=Q_{1-\alpha}(\bm{w}^{*})$.
-Since $\bm{w}^{*}$ and $q^{*}$ are deterministic once $\alpha$ and the true distribution are given, We can briefly denote the optimal prediction set for $x$ by
-$\widehat{C}^{*}_{1-\alpha}(x):=\widehat{C}_{1-\alpha}(x,\bm{w}^{*},q^{*}).$
-From a theoretical perspective, the expected size of the prediction set (at the population level) obtained using $\widehat{\bm{w}}$ cannot be smaller than the one provided by $\bm{w}^{*}$.
-In the subsequent analysis, we will investigate the difference between the prediction set sizes yielded by $\widehat{\bm{w}}$ and $\bm{w}^{*}$.
-To facilitate this analysis, we will introduce and rely on the following events.
-We define
-$\Omega(\mathcal{I},\eta)=\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq q\}-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq q\}]\right|\leq\eta$ (7)
-and
-$\Gamma(\mathcal{I},\xi)=\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(X_{i},y)\rangle\geq q\}-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,y)\rangle\geq q\}]\right|\leq\xi.$ (8)
+$\bm{w}^{*}\in\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}\left[|\widehat{C}_{1-\alpha}(X,\bm{w},Q_{1-\alpha}(\bm{w}))|\right].$
+We define $q^{*}=Q_{1-\alpha}(\bm{w}^{*})$.
+Since $\bm{w}^{*}$ and $q^{*}$ are deterministic once $\alpha$ and the true distribution are given, we can succinctly denote the optimal prediction set for $x$:
+$\widehat{C}_{1-\alpha}^{*}(x):=\widehat{C}_{1-\alpha}(x,\bm{w}^{*},q^{*}).$
+Our analysis of efficiency focuses on the difference between the size of the prediction set produced by the proposed methods using $\widehat{\bm{w}}$ and the expected size of $\widehat{C}^{*}_{1-\alpha}(x)$.
+In other words, we analyze the discrepancy between the prediction set sizes generated by $\widehat{\bm{w}}$ and the optimal $\bm{w}^{*}$.
+We will demonstrate that this difference is negligible given a sufficiently large dataset.
+We note that this analysis differs from verifying whether $\widehat{\bm{w}}$ converges to $\bm{w}^{*}$, which would require assumptions about the smoothness of the function mapping the weight to the prediction set size.
+Instead, our analysis makes fewer assumptions on the optimal prediction set size.
+3.2 Results by Vapnik–Chervonenkis Theory
+We begin by defining the following events on the probability space of pairs of exchangeable random variables $(X_{i},Y_{i})_{i\in\mathcal{I}}$. The event $\Omega(\mathcal{I},\eta)$ is defined as:
+$\begin{split}\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}&\Big{|}\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq q\}\\ &-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq q\}]\Big{|}\leq\eta,\end{split}$ (7)
+where $(X,Y)$ has the same distribution as $(X_{1},Y_{1})$. $\Gamma(\mathcal{I},\xi)$ denotes the event
+$\begin{split}\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\Big{|}&\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(X_{i},y)\rangle\geq q\}\\ &-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,y)\rangle\geq q\}]\Big{|}\leq\xi,\end{split}$ (8)
+where we recall that $K$ is the number of classes of the classification task.
 $\Omega(\mathcal{I},\eta)$ is about the uniform concentration of the coverage, and $\Gamma(\mathcal{I},\xi)$ is about the uniform concentration of the prediction set size.
@@
 (5),
-$\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\mathbf{1}\big{\{}y\in\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{\}}\quad{\text{and}}\quad\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\big{|}\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{|}.$
+[…]
+$\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\big{|}\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{|}.$
 The following lemma shows that these two events hold with high probability.
@@
 Suppose the samples in $\mathcal{I}$ are i.i.d., then
-$\mathbb{P}\left(\Omega\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (9)
-and
-$\mathbb{P}\left(\Gamma\left(\mathcal{I},8K\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+K\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (10)
-This lemma says that $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with high probability if $\eta,\xi\gtrsim\sqrt{\frac{d\log|\mathcal{I}|}{|\mathcal{I}|}}$.
-Now we are ready to introduce the statistical guarantees of the methods.
+  • (a)
+    $\Omega\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)$ hold with probability at least $1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right)$.
+  • (b)
+    $\Gamma\left(\mathcal{I},8K\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+K\delta\right)$ hold with probability at least $1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right)$.
+The proof of this lemma appears in Section B in the appendix.
+This lemma establishes that $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with high probability provided that $\eta,\xi\gtrsim\sqrt{\frac{d\log|\mathcal{I}|}{|\mathcal{I}|}}$.
+The proof, presented in the appendix, applies subgraph classes from Vapnik-Chervonenkis (VC) theory.
+Additionally, if the inequalities in Equations Eq.
+(7) and Eq.
+(8) hold for $\bm{w}\in\mathbb{R}^{d}$, they naturally extend to $\bm{w}\in\mathcal{W}\subset\mathbb{R}^{d}$.
+Remark.
+Part (b) of Lemma 1 limits the generalization of our current theoretical results to classification tasks.
+Compared to the bound in Equation Eq.
+(7), the bound in Equation Eq.
+(8) involves a summation over $[K]$.
+In the proof, this is handled using a union bound.
+However, for regression tasks where prediction set size is typically measured by the Lebesgue measure, a union bound over $K$ is not directly applicable.
+Therefore, our current work focuses exclusively on classification tasks.
+Extending the proposed method to regression tasks is left for future research.
 3.3 Consistency of VFCP
@@
 Theorem 1.
-Let $\widehat{C}_{1-\alpha}^{\text{VFCP}}(x)$ be the output of Algorithm 2 with the setting of VFCP, and modified as mention above.
+Let $\widehat{C}_{1-\alpha}^{\text{VFCP}}(x)$ be the output of Algorithm 2 with the setting of VFCP, and modified as mentioned above.
 Suppose the samples in $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable, then for $(X,Y)$ in the test set, the coverage probability
-$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\mid\widehat{\bm{w}}\right)\geq 1-\alpha.$
-Moreover, if $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, and $\alpha^{\prime}+\eta_{1}+\eta_{3}\leq\alpha$, then for any $X$ in the test set,
-$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\big{|}\mid\widehat{\bm{w}},\widehat{Q}_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\ \right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{2},$
+$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\right)\geq 1-\alpha.$
+Moreover, if $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, and $\alpha^{\prime}+\eta_{1}+\eta_{3}\leq\alpha$, then for test sample $X$,
+$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\big{|}\right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{2},$
 where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha^{\prime})\rceil-\eta_{3}.$
-In the setting of EFCP, $\mathcal{I}_{1}=\mathcal{I}_{2}$.
-To emphasize their roles in the algorithm, we use different notation in the theorem.
-$\widehat{\bm{w}}$ and $\widehat{Q}_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ depend on the dataset $\mathcal{I}_{\textrm{train}}.$ As long as $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, then the coverage and the expected prediction set size satisfies the inequalities in the theorem.
+The proofs of all theorems appear in Section C and Section D and the appendix.
+The validity of VFCP requires minimal assumption (exchangeability), while it is less efficient than other method. We will verify this fact in empirical studies.
 3.4 Consistency of EFCP
@@
 Theorem 2.
-Let $\widehat{C}_{1-\alpha}^{\text{EFCP}}(x,\widehat{\bm{w}})$ be the output of Algorithm 2 with the setting of EFCP.
+Let $\widehat{C}_{1-\alpha}^{\text{EFCP}}(x)$ be the output of Algorithm 2 with the setting of EFCP.
 Suppose the data in $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ are independent and $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ is satisfied.
 Then for $(X,Y)$ in the test set, the coverage probability
-$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{EFCP}}(X;\widehat{\bm{w}})\mid\widehat{\bm{w}}\ \right)\geq 1-\alpha_{1}\quad\text{where}\quad 1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
+$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{EFCP}}(X)\right)\geq 1-\alpha_{1}\quad\text{where}$
+11 11 $-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
 Moreover, if $\Gamma(\mathcal{I}_{\textrm{train}},\xi_{\text{train}})$ is satisfied, then for any $X$ in the test set,
-$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{EFCP}}(X,\widehat{\bm{w}})\big{|}\mid\widehat{\bm{w}}\ \right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{\text{train}}.$
+$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{EFCP}}(X)\big{|}\right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{\text{train}}.$
 3.5 Consistency of DLCP
@@
 In this case, the test set and $\widehat{\bm{w}}$ become dependent.
-We will present the result under a high probability event about the prediction set size.
+The prediction set size is presented differently than previous results.
 Theorem 3.
-Let $\widehat{C}_{1-\alpha}^{\text{DLCP}}(x))$ be the output of Algorithm 2 with the setting of DLCP.
+Let $\widehat{C}_{1-\alpha}^{\text{DLCP}}(x))$ be the output of Algorithm 2 for a test sample $x$ with the setting of DLCP.
 Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
-$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})\}\geq 1-\alpha_{1}-\eta_{\text{test}}\quad\text{where}\quad 1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
+$\frac{1}{|\mathcal{I}_{\textrm{test}}|}$ $\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})\}\geq 1-\alpha_{1}-\eta_{\text{test}}\quad\text{where}$
+$1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
 In addition, if $\Gamma(\mathcal{I}_{\textrm{test}},\xi_{\text{test}})$ holds, then
@@
 Similar as the DLCP setting, $\widehat{\bm{w}}$ depends on the test set, so the result of prediction set size is presented in a similar way.
+In the following theorem, we use $\mathcal{I}_{2}$ to denote $\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$.
 Theorem 4.
 Let $\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x,\widehat{\bm{w}}))$ be the output of Algorithm 2 with the setting of DLCP+.
-Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
-$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x_{i},\widehat{\bm{w}})\}\geq 1-\alpha^{\prime}-\eta_{\text{test}}\quad\text{where}\quad 1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
-In addition, if $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for $X$ in the test set,
+Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion
+$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x_{i},\widehat{\bm{w}})\}\geq 1-\alpha^{\prime}-\eta_{\text{test}}$
+$\text{where}\quad 1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
+In addition, if $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for test sample $X$,
 $\mathbb{E}[|\widehat{C}_{1-\alpha}^{\text{DLCP+}}(X)|]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha^{\prime}}(X)\big{|}\right]+2\xi_{2}.$
-We note that in the setting of DLCP+, $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. We keep use $\mathcal{I}_{2}$ in the theorem to avoid long notation.
+3.7 Conclusion of Theoretical Results
+We integrate the results of Lemma 1 with our main theorems.
+The Lemma 1 indicates that as the size of the set $\mathcal{I}$ approaches infinity, the events $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with probability $1-o(1)$ even if $\eta,\xi\to 0$ is at a sufficient slow rate.
+Consequently, the terms $\eta_{i}$ and $\xi_{i}$ in our theorems become negligible.
+This suggests that the proposed methods achieve coverage rates close to $1-\alpha$ and exhibit near-optimal efficiency.
 4 Experiments
-We conducted an experiment to compare the performance of various score functions and their weighted combinations for prediction set evaluation.
-We focused on three powerful score functions: THR (Sadinle et al., 2019), APS (Romano et al., 2020), and RANK (Luo & Zhou, 2024).
-Throughout the experiment, we assumed that a pretrained classifier $\widehat{p}$ was available, and we focused on the split of the dataset unseen during the training of the pretrained classifier [^1] The unseen dataset was further divided into $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$.
-The naming for $\mathcal{I}_{\textrm{train}}$ is not to be confused with the dataset for training.
-.
-The three basic score functions used in our experiment were:
-Least Ambiguous Set Values Classifier (THR) (Sadinle et al., 2019).
-The score function of THR is defined as:
-$s_{\text{THR}}(x,y)=\widehat{p}_{y}(x).$
-This score function is straightforward: it assigns a higher score to labels with a higher estimated probability.
-The prediction set includes labels with the highest estimated probabilities.
-This is the score function we have to include because if $\widehat{p}$ is the true posterior probability, then the score function itself can achieve the smallest expected prediction set size.
-Adaptive Prediction Set (APS) (Romano et al., 2020).
-The score function of APS is defined as:
-$s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\},$
-where $\mathbf{1}\{\cdot\}$ is the indicator function.
-This score function can be interpreted as the complement of the $p$-value for label $k$.
-It measures the sum of the estimated probabilities of all labels that have the same or a smaller estimated probability than label $k$.
-Rank-based Score Function (RANK) (Luo & Zhou, 2024).
-The score function of RANK is defined as:
-$s_{\text{RANK}}(x,y)=\frac{|\{k^{\prime}\in[K]:\widehat{p}_{k^{\prime}}(x)<\widehat{p}_{k}(x)\}|}{K-1},$
-This score function assigns a score based on the rank of the estimated probability $\widehat{p}_{y}(x)$ among all the estimated probabilities for input $x$.
-The rank is divided by $K-1$ so that the range of the score is from 0 to 1.
-The prediction set gives higher priority to labels with larger ranks.
-The three score functions, THR, APS, and RANK, employ different strategies for assigning scores based on the estimated probabilities.
-THR directly utilizes the estimated probabilities as scores.
-APS, on the other hand, considers the cumulative probability of labels that have the same or lower estimated probabilities compared to the label of interest.
-RANK, in contrast, focuses on the relative ranking of the estimated probabilities among all possible labels.
-It is important to note that all these score functions preserve the order of the labels, which means that, for fixed $x$, the order of labels based on their estimated probabilities, $\widehat{p}_{y}(x)$, remains the same when ranked according to their scores, $s(x,y)$.
-Figure 2: Boxplot comparison of different score functions at a significance level of $\alpha=0.01$.
-Our weighted combination method achieves the guaranteed coverage of 99% while maintaining the smallest prediction set size among all the compared score functions.
-Experiments were conducted on the CIFAR-100 dataset (Alex, 2009) using a pretrained RepVGG model (Ding et al., 2021).
-Our experiments focus on this dataset because of the following facts we found in our experiments.
-  1. 1.
-    This dataset has a well pretrained model, RepVGG (Ding et al., 2021) as our benchmark model.
-  2. 2.
-    This is one of the most challenging classification tasks in computer vision because of its large number of classes.
-  3. 3.
-    There is no existing score functions proposed in earlier work that obviously outperforms all others.
-Given these three aspects, it will be easier for us to present the effectiveness of the weighted score.
-The 10000 testing images, which were not used during the pretraining of the model, were used as the $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ sets.
-The experiments were performed for different significance levels $\alpha$ ranging from 0.01 to 0.05 and two different ratios of $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$, namely 80:20 and 99:1.
-For each combination of $\alpha$ and data split ratio, 100 runs with different index splits were conducted to ensure robustness.
-The main focus of the experiments was to compare the performance of the proposed weighted score function with three constituent base score functions: APS, THR, and RANK.
-Additionally, the proposed method was compared with two competitive baseline score functions, RAPS (Angelopoulos et al., 2021) and SAPS (Huang et al., 2024).
-Figure 2 displays the boxplot of the coverage and size of our proposed method alongside five other score functions at a significance level of $\alpha=0.01$.
-To further analyze the performance across different significance levels, Figure 3 presents a plot of the size against coverage for the various score functions, with $\alpha$ values ranging from 0.01 to 0.05.
-To ensure a fair comparison and guarantee the coverage property, the first data split method, VFCP, was employed for all methods.
-Our method achieves the guaranteed coverage with the smallest prediction set size among all compared score functions, demonstrating its ability to optimize efficiency while maintaining the desired coverage level.
+4.1 Score Weighting
+We conducted an experiment to compare the performance of various score functions score functions discussed in Section 2.2: THR Sadinle et al.
+(2019), APS Romano et al.
+(2020), and RANK Luo and Zhou (2024) given by a pretrained classifier $\widehat{p}_{y}(x)$.
+Additional details about implementation can be found in Section E in the appendix.
+Throughout the experiment, we assumed that a pretrained classifier $\widehat{p}_{y}(x)$ was available, and the split of the dataset unseen during the training of the pretrained classifier.
+Figure 2: Boxplot comparison of different score functions at a significance level of $\alpha=0.01$ on CIFAR-100.
+Our weighted combination method achieves the guaranteed coverage of 99% while maintaining the smallest prediction set size.
+In the experiments on CIFAR-10 and CIFAR-100, testing images, which were not used during the pretraining of the model, were used as the $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ sets.
+The experiments were performed for different significance levels $\alpha$ ranging from 0.01 to 0.05.
+100 runs with different index splits were conducted to ensure robustness.
+We have additional experiments on data splitting ratio, which can be found in Section F in the supplimentary file.
 Figure 3: Comparison of size vs.
-coverage for different score functions as well as our proposed method across $\alpha$ values ranging from 0.01 to 0.05.
-The results indicate that our weighted combination method (shown in red) consistently outperforms the three constituent basic score functions (APS, THR, RANK) as well as other competitive baseline score functions (RAPS, SAPS) in terms of achieving coverage with smaller prediction set sizes.
-To investigate the impact of various data split methods (Section 2.4), we further compare the performance of our weighted score function under different split approaches.
-Additionally, we include the results of the five individual score functions using the VFCP split method.
-Table 1 offers a comprehensive overview of the coverage and size results at significance levels of $\alpha=0.01$ and 0.05.
-These findings demonstrate the effectiveness of the different split strategies.
+coverage for various score functions and our proposed method across $\alpha$ values (0.01-0.05).
+Our weighted combination method (red) consistently outperforms the other baseline methods by achieving the desired coverage rate with smaller prediction set sizes.
 $\alpha=0.01$ $\alpha=0.05$
 Method Coverage Size Coverage Size
-Weighted (VFCP) 0.990 (0.003) 13.782 (1.114) 0.950 (0.005) 3.890 (0.266)
-Weighted (EFCP) 0.989 (0.003) 13.306 (0.464) 0.949 (0.005) 3.754 (0.096)
-Weighted (DLCP) 0.989 (0.003) 13.298 (0.461) 0.949 (0.005) 3.752 (0.097)
-Weighted (DLCP+) 0.989 (0.003) 13.299 (0.459) 0.949 (0.005) 3.753 (0.097)
+VFCP 0.990 (0.003) 13.782 (1.114) 0.950 (0.005) 3.890 (0.266)
+EFCP 0.989 (0.003) 13.306 (0.464) 0.949 (0.005) 3.754 (0.096)
+DLCP 0.989 (0.003) 13.298 (0.461) 0.949 (0.005) 3.752 (0.097)
+DLCP+ 0.989 (0.003) 13.299 (0.459) 0.949 (0.005) 3.753 (0.097)
 APS 0.990 (0.003) 40.217 (1.786) 0.949 (0.006) 19.545 (1.022)
@@
 SAPS 0.990 (0.003) 25.096 (2.298) 0.950 (0.006) 5.036 (0.302)
-Table 1: Coverage and size of different methods for $\alpha=0.01$ and $\alpha=0.05$ on CIFAR-100 dataset. Results are shown as mean (standard deviation).
-(a) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 80:20.
-EFCP, DLCP, and DLCP+ exhibit similar size results, but DLCP has the smallest coverage and the largest gap from the desired coverage level of $0.951-\alpha=$ 0.95.
-VFCP attains the desired coverage at the cost of having the largest prediction set size.
-(b) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 99:1.
-DLCP achieves the smallest size among the compared methods.
-VFCP attains the desired coverage at the cost of having the largest prediction set size.
-Figure 4: Comparison of coverage and size for different data split methods at $\alpha$=0.05.
+Table 1: Coverage and size of different methods for $\alpha=0.01$ and $\alpha=0.05$ on CIFAR-100 dataset.
+Results are shown as mean (standard deviation).
+The first four methods corresponds to the splitting methods in Section 2.5.
+The primary objective of our experiments was to evaluate the performance of the proposed weighted score function in comparison with three foundational base score functions: APS, THR, and RANK.
+Additionally, we compared our method against two competitive baseline score functions, RAPS Angelopoulos et al.
+(2021) and SAPS Huang et al.
+(2024).
+Figure 2 compares coverage and prediction set sizes across methods at $\alpha=0.01$ on CIFAR-100, while Figure 3 evaluates performance across $\alpha\in[0.01,0.05]$.
+Using VFCP splits for consistency, our method achieves guaranteed coverage with the smallest prediction sets overall.
+The advantages are particularly pronounced on CIFAR-10 across all $\alpha$ values and on CIFAR-100 for $\alpha\leq 0.02$.
+Notably, THR exhibits strong performance on less challenging tasks (CIFAR-10), where our weighting scheme naturally assigns it dominant weights.
+In these cases, score averaging provides limited improvement.
+However, for complex multi-class scenarios (CIFAR-100) or stringent significance levels ($\alpha<0.05$), our method demonstrates clear superiority by optimally combining score functions.
+Furthermore, to explore the influence of different data splitting strategies (as discussed in Section 2.4), we conducted additional comparisons of our weighted score function using various split approaches.
+We also included the performance results of the five individual score functions when utilizing the VFCP split method.
+Table 1 provides a comprehensive summary of the coverage and size metrics at significance levels of $\alpha=0.01$ and $\alpha=0.05$.
+These results highlight the effectiveness of the different data splitting strategies.
+It is important to note that while APS aims to achieve conditional coverage, the other methods, including our approach, do not specifically target conditional coverage.
+Consequently, it is not surprising that the alternative methods exhibit better efficiency compared to APS.
+4.2 Model Weighting
+In addition to choosing the score functions in Section 2.2, we further conducted an experiment to compare the performance of various models and their weighted combinations for prediction set construction on CIFAR-10 and CIFAR-100.
+For CIFAR-10, we used the models ResNet-56, ShuffleNetV2 (1.0x), and VGG16-BN.
+For CIFAR-100, we used the models VGG16-BN, RepVGG-A2, and MobileNetV2 (1.0x).
+The performance of the weighted combination method was compared against individual models.
+Specifically, for each of the four score functions: THR, APS, RAPS, and SAPS, we generated prediction sets by combining the scores of different models using weights selected by our proposed method.
+The results are visualized in Figures 4 and 5, which demonstrate the effectiveness of our weighted combination method in terms of weighing scores of different models.
+Figure 4: Across various score functions, our weighted combination of models outperformed any individual model and achieved optimal size on the CIFAR-10 dataset across $\alpha$ values (0.01–0.05).
+Figure 5: Across various score functions, our weighted combination of models outperformed any individual model and achieved optimal size on the CIFAR-100 dataset across $\alpha$ values (0.01–0.05).
 5 Related Work
-We would like to underscore the connections between our work and existing studies on conformal prediction, particularly those that delve into model aggregation and calibration.
-Regarding model aggregation, Yang & Kuchibhotla (2024) introduce two selection algorithms aimed at minimizing the width of the prediction interval by aggregating and selecting from regression estimators.
-Stutz et al.
-(2023) further extends conformal prediction to scenarios where the ground truth label is ambiguous, employing predictions from multiple labels to construct valid prediction sets.
-In the context of model calibration, temperature scaling (Guo et al., 2017) is utilized as a post hoc method to adjust classifier confidence levels.
-Additionally, Rahimi et al.
-(2020) endeavors to calibrate predicted probabilities from deep learning models while maintaining the rank order of these probabilities.
-Dabah & Tirer (2024); Xi et al.
-(2024) empirically demonstrate that applying post hoc calibration prior to conformal prediction can reduce model overconfidence but tends to increase the size of the prediction interval.
-Lin et al.
-(2021) addresses calibration in regression settings, leveraging kernel regression and localized conformal prediction based on a pretrained neural network mean estimator to create valid and discriminative prediction intervals.
-However, the original split conformal method (Papadopoulos et al., 2002) only utilizes a portion of the data for prediction, leading to reduced data efficiency.
-To address this limitation, several approaches have been proposed to enhance data utilization while maintaining computational efficiency.
-Aggregation methods have emerged as a popular approach to improve data utilization as well as prediction efficiency in conformal prediction.
-Carlsson et al.
-(2014) extended inductive conformal prediction (ICP) by aggregating predictions from multiple ICPs, reducing variance and improving efficiency, albeit more conservatively than ICP.
-The cross-conformal prediction Vovk (2015); Vovk et al.
-(2018) presents a hybrid of inductive conformal prediction and cross-validation, offers computational efficiency comparable to ICP while producing valid predictions empirically.
-The validity of aggregated conformal predictors is not fully understood and requires additional assumptions; a revised definition is proposed Linusson et al.
-(2017) to achieve approximate validity conditional on the underlying learning algorithm’s properties.
-Johansson et al.
-(2014) demonstrated that random forests as the underlying model for conformal regression outperform others in terms of efficiency, using out-of-bag estimates instead of a separate calibration set.
-BARBER et al.
-(2021) introduced the jackknife+ method, which employs leave-one-out cross-validation to compute nonconformity scores for each example in the calibration set, enabling the use of all available data for both training and calibration.
-Kim et al.
-(2020) proposed a similar approach using K-fold cross-validation, where the model is trained on multiple folds of the data, and the predictions from each fold are combined to create a valid prediction set.
-These methods strike a balance between data efficiency and computational complexity by reusing data for both training and calibration.
-Recently, conformal prediction has been extended to handle distribution shifts Tibshirani et al.
-(2019); Gibbs et al.
-(2023); Prinster et al.
-(2024), relaxing the assumption of exchangeability.
-This advancement allows for the application of conformal prediction to a wider range of real-world scenarios, such as data with dependencies Chernozhukov et al.
-(2021); Luo et al.
-(2023); Nettasinghe et al.
-(2023); Luo & Colombo (2024).
-It is worth noting that there is an inherent connection between model aggregation and our approach of weighting non-conformity scores, even though our method does not directly address the scenario of combining multiple models to find an optimal aggregation.
-To illustrate this connection, consider a scenario with three classifiers, including an oracle classifier making perfect predictions.
-A model aggregation method such as boosting (Mukherjee & Schapire, 2010) would assign a weight of 1 to the oracle classifier, disregarding the others.
-Similarly, when weighting non-conformity scores, even without a perfect classifier, the resulting conformal prediction sets will be equivalent to applying the THR score to the oracle classifier.
-The optimally chosen weighted non-conformity score can identify and prioritize the most informative and reliable scores, mimicking an oracle classifier.
-This equivalence highlights our method’s potential to achieve optimal performance by leveraging multiple non-conformity scores, even without a perfect classifier, approximating the performance of an ideal classifier through appropriate weight assignment.
-6 Conclusion
-In conclusion, our weighted aggregation approach addresses the challenge of constructing efficient and valid prediction sets in multi-class classification by learning optimal weights for combining different score functions using various data split procedures.
-This adaptability allows our method to adjust to the specific characteristics of the dataset, potentially outperforming methods that rely on a single score function.
-Our theoretical analysis establishes finite-sample guarantees for coverage, ensures the validity of the prediction sets, and derives oracle inequalities that quantify the expected set size of our approach relative to an optimal weight.
-Experiments on the CIFAR-100 dataset demonstrate the superior performance of our weighted approach compared to single-score methods, consistently achieving the guaranteed coverage while maintaining the smallest prediction set sizes, which highlights the potential of our method to achieve optimal performance by leveraging the strengths of multiple non-conformity scores.
-Future research directions include exploring the application of our approach to other domains and investigating the integration of advanced machine learning techniques to further enhance the performance and efficiency of conformal prediction sets.
-Appendix A Proof of Lemma 1
-A.1 VC Dimension of the Subgraph Classes
+Our work builds upon existing research in conformal prediction, particularly focusing on model averaging and calibration techniques.
+In the realm of model aggregation, Yang et al.
+Yang and Kuchibhotla (2024) introduced two selection algorithms designed to minimize the width of prediction intervals by aggregating and selecting from multiple regression estimators.
+The technique of VC dimension has been applied in Yang and Kuchibhotla (2024) and Candès et al.
+(2023).
+In Section D.2 of Yang and Kuchibhotla (2024), the uniform probability bound is over a finite set.
+In Candès et al.
+(2023), the authors considers maximization over real numbers.
+They have not consider the techiniques subgraph theory like our approach.
+The methods proposed in Carlsson et al.
+(2014); Linusson et al.
+(2017) can aggregate different prediction models to perform conformal prediction, but the methods do not focus on optimizing the efficiency.
+A concurrent preprint Liang et al.
+(2024) examines parametrized score functions $s_{\lambda}(x,y)$ for $\lambda\in\Lambda$ in regression settings.
+While their general framework handles arbitrary parameter spaces, controlling the Rademacher complexity becomes challenging except in specialized cases.
+Our approach differs fundamentally by combining pre-defined score functions linearly.
+This structural constraint enables theoretical guarantees dependent only on the number of constituent scores rather than the complexity of function spaces.
+6 Conclusion and Discussion
+Our weighted score aggregation method enables efficient and valid prediction set construction for multi-class classification through optimized combinations of score functions and strategic data splitting.
+Theoretically, we establish finite-sample coverage guarantees and oracle inequalities quantifying the efficiency gap between our method and optimal weights.
+Empirically, experiments demonstrate consistent maintenance of coverage requirements with minimal prediction set sizes compared to single-score baselines.
+This work bridges model averaging and conformal prediction, providing a flexible framework for uncertainty quantification that adapts to dataset characteristics through optimal score combinations.
+These results suggest several directions for future work:
+  1. 1.
+    Regression Extension: Algorithmic adaptation to regression requires new theoretical tools due to continuous output spaces, where finite-class union bounds become inapplicable.
+Potential approaches include metric entropy analysis or covering number techniques.
+  2. 2.
+    Optimization Enhancement: Developing gradient-based alternatives to grid search, such as differentiable conformal objectives or online weight adaptation during model training, would improve scalability with many score functions.
+Appendix A Grid Search Implementation
+To solve Eq. (4), we discretize the probability simplex:
+$\Delta^{d-1}=\left\{\bm{w}\in\mathbb{R}_{+}^{d}:\sum_{j=1}^{d}w_{j}=1\right\}$
+using a grid resolution $\varepsilon=0.01$. Candidate weights are generated as:
+$\mathcal{W}=\left\{\bm{w}=(k_{1}\varepsilon,\ldots,k_{d}\varepsilon)\,\,\bigg{|}\,\,k_{j}\in\mathbb{N},\,\,\sum_{j=1}^{d}k_{j}=\lfloor 1/\varepsilon\rfloor\right\},$
+yielding $\binom{\lfloor 1/\varepsilon\rfloor+d-1}{d-1}$ distinct weight vectors.
+For $d=3$ scores, this produces 5,151 candidates, ensuring comprehensive coverage of the parameter space while remaining computationally tractable through parallel evaluation.
+Appendix B Proof of Lemma 1
+B.1 VC Dimension of the Subgraph Classes
 Proposition 1.
@@
 ∎
-A.2 Proof of the Lemma
+B.2 Proof of the Lemma
 This proof basically follows from Vapnik–Chervonenkis theory.
 We will use the theorems and lemmas in Wainwright (2019) as reference.
-The proof of Eq.
-(9) and Eq.
-(10) are almost identical.
-We will focus on Eq.
-(9).
+The proof of both parts of the lemma are almost identical.
+We will focus on part (a).
 Let $\epsilon_{i},i\in\mathcal{I}$ be i.i.d.
@@
 $\mathcal{R}(\mathcal{I})\leq 4\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}$ (12)
-We combine Eq. (13) and Eq. (14) to obtain Eq. (9). The proof of Eq. (10) follows analogously, using the VC dimension result for the function class in Eq. (12). Let us define
+We combine Eq.
+(11) and Eq.
+(12) to obtain the result of part (a).
+The proof of part (b) of the lemma follows analogously, using the VC dimension result for the function class in Eq.
+(10).
+Let us define
 $\Gamma_{y}(\mathcal{I},\eta)=\sup_{w\in\mathbb{R}^{d},t\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle w,s(X_{i},y)\rangle\geq t\}-\mathbb{E}[\mathbf{1}\{\langle w,s(X,y)\rangle\geq t\}]\right|\leq\eta.$
@@
 $\bigcup_{y\in[K]}\Gamma_{y}(\mathcal{I},\eta)\subseteq\Gamma(\mathcal{I},K\eta).$
-Taking the union bound on $y\in[K]$ implies Eq. (10).
-A.3 Bounds for Coverage Probability
+Taking the union bound on $y\in[K]$ implies the result of part (b).
+Appendix C Preliminary Lemmas for the Theorems
+C.1 Bounds for Coverage Probability
 The following result requires the data for optimizing $\widehat{\bm{w}}$ are independent with the calibration set and the test set.
@@
 The remaining proof has similar argument as Eq. (13) and Eq. (14), and is omitted here. ∎
-A.4 Bounds for Prediction Set Size
+C.2 Bounds for Prediction Set Size
 Lemma 5.
@@
 The proof is complete. ∎
-Appendix B Proof of the Theorems
-B.1 Proof of Theorem 1
+Appendix D Proof of the Theorems
+D.1 Proof of Theorem 1
 Under the assumption of the theorem, the setting of VFCP satisfies the condition of Lemma 2.
@@
 This implies the coverage probability of the theorem.
-B.2 Proof of Theorem 2
+D.2 Proof of Theorem 2
 Under the assumption of the theorem, the setting of VFCP satisfies the condition of Lemma 3.
@@
 This implies the coverage probability of the theorem.
-B.3 Proof of Theorem 3
+D.3 Proof of Theorem 3
 Since $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$, one can verify that the theorem is the direct result of Lemma 4 and Lemma 5.
-B.4 Proof of Theorem 4
+D.4 Proof of Theorem 4
 Similar as the setting of DLCP, the coverage probability is the direct result of Lemma 4. For the prediction set size, suppose $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for $X$ in the test set,
@@
 $\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*})\}\right]+2\xi_{2}.$
+Appendix E Additional Discussion on Score Functions
+The three basic score functions used in our experiment were:
+Least Ambiguous Set Values Classifier (THR) Sadinle et al. (2019).
+The score function of THR is defined as:
+$s_{\text{THR}}(x,y)=\widehat{p}_{y}(x).$
+This score function is straightforward: it assigns a higher score to labels with a higher estimated probability.
+The prediction set includes labels with the highest estimated probabilities.
+This is the score function we have to include because if $\widehat{p}$ is the true posterior probability, then the score function itself can achieve the smallest expected prediction set size.
+Adaptive Prediction Set (APS) Romano et al. (2020).
+The score function of APS is defined as:
+$s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\},$
+where $\mathbf{1}\{\cdot\}$ is the indicator function.
+This score function can be interpreted as the complement of the $p$-value for label $k$.
+It measures the sum of the estimated probabilities of all labels that have the same or a smaller estimated probability than label $k$.
+Rank-based Score Function (RANK) Luo and Zhou (2024).
+The score function of RANK is defined as:
+$s_{\text{RANK}}(x,y)=\frac{|\{k^{\prime}\in[K]:\widehat{p}_{k^{\prime}}(x)<\widehat{p}_{k}(x)\}|}{K-1},$
+This score function assigns a score based on the rank of the estimated probability $\widehat{p}_{y}(x)$ among all the estimated probabilities for input $x$.
+The rank is divided by $K-1$ so that the range of the score is from 0 to 1.
+The prediction set gives higher priority to labels with larger ranks.
+The three score functions, THR, APS, and RANK, employ different strategies for assigning scores based on the estimated probabilities.
+THR directly utilizes the estimated probabilities as scores.
+APS, on the other hand, considers the cumulative probability of labels that have the same or lower estimated probabilities compared to the label of interest.
+RANK, in contrast, focuses on the relative ranking of the estimated probabilities among all possible labels.
+It is important to note that all these score functions preserve the order of the labels, which means that, for fixed $x$, the order of labels based on their estimated probabilities, $\widehat{p}_{y}(x)$, remains the same when ranked according to their scores, $s(x,y)$.
+Appendix F Additional Experiments with Different Data Splitting Ratio
+(a) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 80:20.
+EFCP, DLCP, and DLCP+ exhibit similar size results, but DLCP has the smallest coverage and the largest gap from the desired coverage level of $0.951-\alpha=$ 0.95.
+VFCP attains the desired coverage at the cost of having the largest prediction set size.
+(b) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 99:1.
+DLCP achieves the smallest size among the compared methods.
+VFCP attains the desired coverage at the cost of having the largest prediction set size.
+Figure 6: Comparison of coverage and size for different data split methods at $\alpha$=0.05.